### v1

# Relative Obstructions and Spectral Diagnostics for Sheaves on Cell Complexes

Abstract
Many structured systems admit locally consistent descriptions that nevertheless fail to globalize when constrained by an ambient reference or feasibility condition. Diagnosing such failures is naturally an evaluative problem: given a fixed model and a grounding, can one determine whether they are structurally compatible, and if not, identify the nature and localization of the obstruction?
In this work, we introduce a sheaf-theoretic and spectral framework for evaluating structural inconsistency as a relative phenomenon. A model is represented by a cellular sheaf $\mathcal{F}$ on a cell complex, together with a morphism into a grounding sheaf $\mathcal{W}$ encoding admissible global behavior. Failure of compatibility is captured by the mapping cone of this morphism, whose cohomology computes the relative groups $H^{*}(K;\mathcal{F},\mathcal{W})$ and separates intrinsic obstructions from inconsistencies induced by the grounding.
Beyond exact cohomological classification, we develop spectral witnesses derived from regular and mapping-cone Laplacians. The spectra of these operators provide computable, quantitative indicators of inconsistency, encoding both robustness and spatial localization through spectral gaps, integrated energies, and eigenmode support. These witnesses enable comparison of distinct inconsistency mechanisms in fixed systems without learning, optimization, or modification of the underlying representation.
The proposed framework is domain-agnostic and applies to a broad class of structured models where feasibility is enforced locally but evaluated globally.
1 Introduction: Structural Inconsistency as a Relative Obstruction
Many structured systems admit locally consistent descriptions that nevertheless fail to globalize. Such failures arise across diverse settings—logical constraint systems, distributed data integration, networked models, and abstract representations—where compatibility is enforced only through local relations. A central mathematical question is therefore not how to generate or repair such structures, but how to evaluate whether a given system is globally feasible relative to a fixed reference.
In this work, we study structural inconsistency as an inverse or verification problem. Given a hypothesized relational structure and an ambient grounding that specifies admissible global behavior, can one certify whether the structure is compatible with the grounding, and if not, identify the nature and localization of the obstruction? Our focus is on diagnosing inconsistency in a fixed system, rather than modifying or learning a new representation.
We formulate this problem using the language of cellular sheaves. A sheaf $\mathcal{F}$ encodes the local behavior of a model on a cell complex $K$, while a grounding sheaf $\mathcal{W}$ represents an ambient reference or feasibility constraint. Compatibility between the two is expressed by a cochain map $\epsilon:\mathcal{F}\to\mathcal{W}$. Failure of sections of $\mathcal{F}$ to extend compatibly into $\mathcal{W}$ is then captured by the mapping cone of $\epsilon$, whose cohomology computes the relative groups $H^{*}(K;\mathcal{F},\mathcal{W})$. This construction distinguishes obstructions intrinsic to the model from those that arise only under grounding, providing a principled algebraic notion of relative inconsistency.
While mapping cones and relative cohomology are classical tools in homological algebra, their use here is not classificatory but diagnostic. Rather than relying solely on exact cohomological invariants, we introduce Laplacian operators associated to both the original sheaf and the mapping cone, and study their spectra as quantitative indicators of structural inconsistency. These spectra encode not only the presence of obstructions, but also their robustness and spatial localization, enabling evaluation beyond binary consistency.
Contributions.
The contributions of this work are as follows.
  1. 1.
    We introduce a relative sheaf-theoretic formulation of structural inconsistency based on the mapping cone of a morphism between a model sheaf and a grounding sheaf. The resulting relative cohomology separates intrinsic topological obstructions from incompatibilities induced by the grounding, and provides a unified algebraic framework for reasoning about feasibility in fixed systems.
  2. 2.
    We develop spectral witnesses derived from regular and mapping-cone Laplacians as computable diagnostics of inconsistency. By analyzing spectral gaps, integrated energies, and eigenmode localization, we obtain global and local indicators that quantify the degree of inconsistency and attribute it to specific cells. These witnesses are stable under bounded perturbations of the harmonic filtration and exhibit robustness to moderate noise in controlled experiments.
  3. 3.
    We emphasize an explicitly evaluative perspective. All sheaves, grounding maps, and operators are constructed from a given model or system; no learning, optimization, or representation adjustment is performed. The resulting spectral diagnostics act as an external evaluation layer, enabling comparison and analysis of distinct inconsistency mechanisms without altering the underlying structure.
Note.
Readers primarily interested in practical diagnostics may safely skip Appendix A.
Note.
Counterfactual feasibility will appear as a motivating example throughout the paper, but the framework applies to general structural compatibility problems beyond causal interpretation.
2 Related Work
Metric and Geometric Approaches to Sheaf Consistency
Related notions of inconsistency have been studied from a geometric perspective via sheaves of pseudometric spaces, where inconsistency is quantified by a sheaf consistency radius measuring the minimal perturbation in a chosen pseudometric required to restore global consistency [13]. Such approaches relax constraints at the level of assignments to a single sheaf, quantifying deviation from exact global sections in a chosen pseudometric. By contrast, the present work adopts an algebraic and explicitly relative viewpoint: inconsistency is diagnosed through the mapping cone of a morphism between two coefficient sheaves—a model sheaf and an ambient grounding—thereby separating intrinsic obstructions from grounding-induced incompatibilities without appealing to metric relaxation.
Ordinary sheaf cohomology for inconsistency.
Ghrist and Cooperband [7] study geometric paradoxes through sheaf-theoretic torsors, with non-trivial first cohomology classes quantifying global inconsistency in a single coefficient sheaf. However, this framework analyzes intrinsic obstructions of a single coefficient sheaf. While mapping cones and relative cohomology are classical tools in homological algebra, they are not employed in this setting to diagnose inconsistency between a model and an ambient grounding, nor are they operationalized as evaluative or spectral diagnostics.
Network sheaves and abstraction morphisms.
Related sheaf-theoretic frameworks for abstraction have been developed by D’Acunto and collaborators [5, 4], who model structured systems and abstraction maps between them using network sheaves and sheaf morphisms. In these works, global sections and Laplacian kernels are used to characterize structural consistency of sections across abstraction levels. The emphasis in this line of work is on achieving or learning absolute consistency of a single network sheaf under abstraction, often by modifying the abstraction map or the underlying representation, rather than on extracting relative invariants associated to a fixed pair of coefficient sheaves. In particular, no mapping cone constructions or relative spectral diagnostics are introduced. The present work complements this perspective by focusing on relative (co)homological obstructions and their spectral signatures, enabling a separation between intrinsic inconsistency and grounding-induced incompatibility.
Spectral methods / Laplacians.
Spectral methods based on graph and sheaf Laplacians provide an analytic bridge between discrete topology and quantitative diagnostics. Classical combinatorial Hodge theory establishes that the kernels of degree-wise Laplacians recover cohomology groups, while spectral gaps encode robustness of these invariants [10]. Building on this foundation, recent work on sheaf Laplacians leverages spectral energy and harmonic sections to analyze global consistency and data fusion in distributed systems [9]. Our approach builds on this tradition but departs from it in two essential respects. First, we introduce relative Laplacians arising from mapping-cone constructions associated to a morphism of coefficient sheaves. Second, we use their spectra explicitly as evaluative witnesses–quantifying and localizing inconsistency in a fixed system–rather than as representations or optimization objectives.
Our Work.
The present work differs from these approaches in three essential ways. First, we model grounding-relative inconsistency (sometimes motivated by counterfactual feasibility) as a relative obstruction, formalized via the mapping cone of a cochain map between a sheaf of model behavior and an ambient grounding sheaf, yielding relative cohomology groups $H^{j}(K;\mathcal{F},\mathcal{W})$. Second, rather than focusing on exact cohomological classification, we introduce spectral witnesses—derived from regular and cone Laplacians—that quantify and localize inconsistency through spectral gaps, integrated energies, and cell-wise eigenmode support. Third, our framework is explicitly evaluative and model-agnostic: the sheaf and its grounding are constructed from a fixed model or system, and the resulting spectral invariants diagnose distinct inconsistency mechanisms without learning or modifying the underlying representation.
While the underlying homological constructions are classical, the contribution of the present work lies in their systematic use as computable, spectral diagnostics for distinguishing and localizing distinct inconsistency mechanisms in a fixed model–grounding pair.
3 Preliminaries
3.1 Graphs, Cellular Sheaves, and Cochains
Let $K$ denote a finite dimensional cell complex, which we identify with a graph consisting of a vertex set $K_{0}$ and an edge set $K_{1}$. We interpret $K$ as a combinatorial domain on which local data and compatibility constraints are defined.
A cellular sheaf $\mathcal{F}$ on $K$ assigns a finite-dimensional vector space $\mathcal{F}(\sigma)$ to each cell $\sigma\in K$, together with linear restriction maps
$\rho_{\sigma\tau}\colon\mathcal{F}(\sigma)\to\mathcal{F}(\tau)\quad\text{whenever }\tau\subseteq\sigma,$
satisfying the usual functorial consistency conditions. In the present work, stalks over vertices encode local state or feature information, while stalks over edges encode pairwise compatibility constraints.
Associated to a cellular sheaf is its cochain complex
$C^{j}(K;\mathcal{F})=\bigoplus_{\sigma\in K_{j}}\mathcal{F}(\sigma),$
equipped with coboundary operators $d^{j}\colon C^{j}\to C^{j+1}$ induced by the restriction maps and oriented incidence structure of $K$. The resulting cohomology groups $H^{j}(K;\mathcal{F})$ measure global consistency properties of the local data. We refer to [3, 6] for general background on cellular sheaves and their cohomology.
3.2 Sheaf Laplacians and Spectral Properties
Given a cellular sheaf $\mathcal{F}$ on $K$, the associated cochain complex admits a family of degree-wise Laplacian operators defined by
$L_{j}=d^{j-1}(d^{j-1})^{\ast}+(d^{j})^{\ast}d^{j},$
where $(\cdot)^{\ast}$ denotes the adjoint with respect to a chosen inner product on each cochain space. Each $L_{j}$ is symmetric and positive semidefinite, and its kernel is canonically isomorphic to the cohomology group $H^{j}(K;\mathcal{F})$.
Remark 1 (Degree-dependent interpretation of obstruction).
We emphasize that the interpretation of sheaf cohomology and Laplacian kernels depends on degree. In degree 0, nonvanishing $H^{0}(K;\mathcal{F})$ corresponds to the existence of a global section, while vanishing indicates the absence of any globally consistent assignment. In contrast, for degrees $j\geq 1$, nonvanishing $H^{j}(K;\mathcal{F})$ represents an obstruction to extending compatible local sections to a global one, whereas vanishing indicates that no such obstruction exists. Throughout this paper, the term obstruction is used in this degree-dependent sense: as failure of global existence in degree 0, and failure of extendability in higher degrees.
This construction extends classical combinatorial Hodge theory from simplicial complexes to cellular sheaves. In particular, the spectrum of $L_{j}$ provides quantitative information beyond the mere presence or absence of cohomology. In degree 0, small nonzero eigenvalues indicate near-failure of global existence, while in higher degrees they indicate near-obstructions to extending locally compatible sections. In both cases, the associated eigenvectors localize deviations from ideal consistency across cells. These degree-aware spectral interpretations form the basis of our consistency analysis via sheaf Laplacians.
The theory of sheaf Laplacians was developed in [9], building on earlier work in combinatorial Hodge theory [10]. We adopt this framework as the foundation for defining spectral witnesses of inconsistency and for extending the analysis to relative and grounded settings.
4 Grounding and Relative Obstructions
To enforce grounding, we construct the mapping cone sheaf $\mathrm{Cone}(\epsilon)$ by adjoining an apex cell encoding the reference stalk. We define the cone sheaf Laplacian $L_{1}^{\mathrm{cone}}$.
4.1 Grounding as a Cochain Map
Definition 1 (Ambient grounding).
Let $K$ be a finite simplicial complex and let $\mathcal{F}$ be a cellular sheaf on $K$. Assume that there exists a fixed finite-dimensional inner-product space $W$ and, for each cell $\sigma\in K$, a linear embedding
$\epsilon_{\sigma}:\mathcal{F}(\sigma)\hookrightarrow W,$
which we refer to as the ambient embedding.
A cochain assignment $x=\{x_{\sigma}\}_{\sigma\in K_{0}}$ is said to be ambiently grounded if there exists a vector
$x_{\ast}\in W$
such that
$x_{\ast}=\epsilon_{\sigma}(x_{\sigma})\qquad\text{for all }\sigma\in K_{0}.$
Equivalently, the local stalk assignments factor through a single global section in the ambient space.
4.2 Mapping Cones and Relative Sheaf Cohomology
We formalize the relationship between two constructions used to enforce global grounding:
  1. ()
    a geometric cone construction obtained by adjoining an apex cell to the base complex, and
  2. ()
    an algebraic mapping cone construction associated to a cochain map into a reference complex.
This clarification is essential for interpreting the resulting Laplacian as a certificate of Grounding-induced inconsistency.
4.2.1 Algebraic mapping cone
Let $K$ be a finite dimensional simplicial complex and let $\mathcal{F}$ be a cellular sheaf on $K$ with cochain complex
$(C^{\ast}(K;\mathcal{F}),d_{\mathcal{F}}).$
Let $W$ be a finite-dimensional inner-product space, and let $\mathcal{W}$ denote the constant sheaf on $K$ with stalk $W$ and identity restriction maps. Its associated cochain complex is denoted
$(C^{\ast}(K;\mathcal{W}),d_{\mathcal{W}}).$
Suppose we are given a family of linear maps
$\epsilon_{\sigma}:\mathcal{F}(\sigma)\longrightarrow\mathcal{W},\qquad\sigma\in K,$
which we regard as an embedding of local stalk data into a global reference space.
If the family $\epsilon={\epsilon_{\sigma}}$ satisfies
$\epsilon_{\tau}\circ d_{\mathcal{F}(\sigma\to\tau)}=d_{\mathcal{W}(\sigma\to\tau)}\circ\epsilon_{\sigma}\qquad\text{for all incidences }\sigma\prec\tau,$ (4.2.1)
then $\epsilon$ defines a sheaf morphism $\mathcal{F}\to\mathcal{W}$ and hence induces a cochain map
$\epsilon^{\ast}:(C^{*}(K;\mathcal{F}),d_{\mathcal{F}})\longrightarrow(C^{\ast}(K;\mathcal{W}),d_{\mathcal{W}}).$
The mapping cone of $\epsilon^{\ast}$ is the cochain complex
$\mathrm{Cone}(\epsilon^{\ast})^{n}:=C^{n+1}(K;\mathcal{F})\;\oplus\;C^{n}(K;\mathcal{W}),$
equipped with differential
$d_{\mathrm{cone}}(\alpha,\beta):=\big(-d_{\mathcal{F}}\alpha,-\epsilon(\alpha)+d_{\mathcal{W}}\beta\big).$
The associated Laplacian, constructed in the usual way from $d_{\mathrm{Cone}}$, will be referred to as the mapping cone Laplacian.
4.2.2 Geometric cone construction
Independently, one may form the geometric cone $\widehat{K}$ of $K$ by adjoining a new apex vertex $\ast$ and the cone cells $\ast\sigma$ for every cell $\sigma\subset K$.
Define a sheaf $\widehat{\mathcal{F}}$ on $\widehat{K}$ by:
  • $\widehat{\mathcal{F}}(\sigma)=\mathcal{F}(\sigma)$ for $\sigma\subset K$,
  • $\widehat{\mathcal{F}}(*)=W$,
  • for each cone incidence $\sigma\prec*\sigma$, the restriction map is $\epsilon_{\sigma}:\mathcal{F}(\sigma)\to W$,
  • all other restrictions are inherited from $\mathcal{F}$ or are identities on $W$.
[…]
4.2.3 Equivalence under compatibility
We now make precise the relationship between the algebraic and geometric constructions.
Proposition 1 (Geometric cone realizes the translated mapping cone).
Let $K$ be a finite simplicial complex and let $\varepsilon:\mathcal{F}\to\mathcal{W}$ be a cellular sheaf morphism, by assuming the compatibility condition 4.2.1, and write $\epsilon^{\ast}:C^{\ast}(K;\mathcal{F})\to C^{\ast}(K;\mathcal{W})$ for the induced cochain map. Let $\widehat{K}$ denote the geometric cone on $K$ with apex vertex $\ast$, and let $\widehat{\mathcal{F}}$ be the cone sheaf on $\widehat{K}$ defined by \$4.2.2.
Then there is a natural isomorphism of cochain complexes
$\bigl(C^{\ast}(\widehat{K};\widehat{\mathcal{F}}),\,d_{\widehat{\mathcal{F}}}\bigr)\;\cong\;\bigl(\mathrm{Cone}(\epsilon^{\ast})[1]^{\ast},\,d_{\mathrm{cone}[1]}\bigr),$
where the translated cone complex is given by the standard shift of the algebraic mapping cone,
$\mathrm{Cone}(\epsilon^{\ast})[1]^{n}\;:=\;\mathrm{Cone}(\epsilon^{\ast})^{n-1}\;=\;C^{n}(K;\mathcal{F})\oplus C^{n-1}(K;\mathcal{W}).$
Under this identification, the translated differential is
$d_{\mathrm{cone}[1]}(x,y)\;:=\;\bigl(d_{\mathcal{F}}x,\ \epsilon^{\ast}x-d_{\mathcal{W}}y\bigr),$
which is precisely the differential induced by the shift convention $d_{A[1]}=-\,d_{A}$ applied to the mapping cone differential
$d_{\mathrm{cone}}(x,y)=\bigl(-d_{\mathcal{F}}x,\ -\epsilon^{\ast}x+d_{\mathcal{W}}y\bigr).$
In particular, with this convention the translated algebraic cone agrees degreewise and differentially with the geometric cone complex.
Proof.
For each $n\geq 0$, every $n$-cell of $\widehat{K}$ is either (i) an original $n$-cell $\tau\in K_{n}$, or (ii) a cone cell $\ast\sigma$ over some (n1)(n-1)-cell $\sigma\in K_{n-1}$.
Hence by definition, we have the canonical isomorphism of vector spaces
$\Phi:C^{n}(\widehat{K};\widehat{\mathcal{F}})\cong\Bigl(\bigoplus_{\tau\in K_{n}}\mathcal{F}(\tau)\Bigr)\ \oplus\ \Bigl(\bigoplus_{\sigma\in K_{n-1}}\mathcal{W}(\sigma)\Bigr)\cong C^{n}(K;\mathcal{F})\ \oplus\ C^{n-1}(K;\mathcal{W})=\mathrm{Cone}(\epsilon^{\ast})[1]^{n},$
where $\Phi$ is defined by
$\Phi^{n}(\alpha)=(\alpha|_{\mathcal{F}},\,(-1)^{n}\alpha|_{\mathcal{W}}).$
To see that $\Phi$ is a cochain isomorphism, we show that $\Phi^{n+1}\circ d_{\widehat{\mathcal{F}}}=d_{\mathrm{cone}[1]}\circ\Phi^{n}$.
Using the apex-first cone orientation from Appendix A.2, a direct inspection of the cellular coboundary yields
[…]
where the alternating signs in $d_{\mathcal{F}}$ and $d_{\mathcal{W}}$ arise from incidence numbers, and the factor $(-1)^{n+1}$ arises solely from the degree shift in the identification $\Phi$.
These contributions assemble to
$\Phi^{n+1}\bigl(d_{\widehat{\mathcal{F}}}\alpha\bigr)=\bigl(d_{\mathcal{F}}x,\ \epsilon^{\ast}x-d_{\mathcal{W}}y\bigr)=d_{\mathrm{cone}[1]}(x,y)=d_{\mathrm{cone}[1]}\bigl(\Phi^{n}(\alpha)\bigr),$
as required. Hence $\Phi$ is an isomorphism of cochain complexes $C^{\ast}(\widehat{K};\widehat{\mathcal{F}})\cong\mathrm{Cone}(\epsilon^{\ast})[1]^{\ast}$. ∎
4.2.4 Failure of compatibility as a defect signal
In practical settings, the compatibility condition 4.2.1 need not hold exactly, especially when restriction maps or embeddings are learned or estimated from data.
Definition 2 (Incidence-level defect).
For any incidence $\sigma\prec\tau$, define the incidence-level defect
$\Delta_{\sigma\to\tau}:=\epsilon_{\tau}\circ d_{\mathcal{F}(\sigma\to\tau)}-d_{\mathcal{W}(\sigma\to\tau)}\circ\epsilon_{\sigma},\qquad\sigma\prec\tau.$
Spectral detection of grounding failure
If $\Delta_{\sigma\to\tau}\neq 0$ for some incidence, then $\epsilon$ does not induce a cochain map and the algebraic and geometric cone constructions no longer coincide. In this case, the residual terms induced by $\Delta_{\sigma\to\tau}$ contribute positive energy to the cone Laplacian. Consequently, the smallest nonzero eigenvalues and associated eigenvectors of the cone Laplacian quantify and localize the failure of global grounding.
Remark 2 (Two regimes).
When $\epsilon$ is a cochain map, the mapping cone ${\rm Cone}(\epsilon)$ is a genuine cochain complex and supports standard derived arguments. In general, we do not assume $\epsilon$ is cochain; the failure $\Delta=\epsilon d_{\mathcal{F}}-d_{\mathcal{W}}\epsilon$ is the incidence-level defect measured by our cone-based spectral diagnostics.
4.2.5 Regular vs. cone Laplacians via relative cohomology.
The regular and cone Laplacians encode complementary cohomological information, and neither subsumes the other. Their relationship is most naturally described through the mapping cone construction and the associated long exact sequence in cohomology.
Lemma 1 (The induced long exact sequence.).
Let $\epsilon:\mathcal{F}\to\mathcal{W}$ be a morphism of cellular sheaves on a finite simplicial complex $K$. Denote by $\mathrm{cone}(\epsilon)$ the mapping cone of $\epsilon$ at the cochain level, the induced long exact sequence of cohomology takes the form
$\cdots\;\longrightarrow\;H^{j}(K;\mathcal{F})\;\xrightarrow{\ (\epsilon_{\ast})^{\ast}\ }\;H^{j}(K;\mathcal{W})\;\longrightarrow\;H^{j}\!\big(K;\mathrm{cone}(\epsilon^{\ast})\big)\;\longrightarrow\;H^{j+1}(K;\mathcal{F})\;\longrightarrow\;\cdots.$
Proof.
See Appendix A.3. ∎
This sequence shows that cone cohomology does not replace regular cohomology; rather, it measures the failure of cohomology classes of $\mathcal{F}$ to be compatible with the ambient reference $\mathcal{W}$.
Note (Cone and relative cohomology).
The sheaf cohomology of the mapping cone is the relative cohomology of the pair of coefficient sheaves $(\mathcal{F},\mathcal{W})$ by definition:
$H^{j}(K;\mathcal{F},\mathcal{W})\;\overset{\rm def}{:=}\;H^{j}\!\big(\mathrm{Cone}(\epsilon)\big)$
Interpretation. The regular Laplacian $L_{j}$ detects nontrivial classes in $H^{j}(K;\mathcal{F})$, corresponding to intrinsic (ungrounded) consistency or obstruction. The cone Laplacian $L_{j}^{\mathrm{cone}}$ detects nontrivial classes in $H^{j}\!\big(\mathrm{Cone}(\epsilon)\big)\cong H^{j}(K;\mathcal{F},\mathcal{W})$, corresponding to classes that fail to admit an ambiently grounded representative. Consequently, comparing the spectra of $L_{j}$ and $L_{j}^{\mathrm{cone}}$ provides a finer spectral witness of inconsistency than either operator alone: it distinguishes between intrinsic cohomological freedom and genuine grounding-induced obstruction.
Remark 3.
In what follows, we work exclusively with the geometric mapping cone, which is sufficient for all diagnostics considered here.
4.3 Interpretation: Lifting and Categorification
This viewpoint is not used technically in what follows, but serves to situate the evaluation problem conceptually.
We interpret the task of diagnosing structural consistency as an inverse problem formulated in categorical terms. Rather than constructing or learning a model, we assume that a relational structure is given in advance, together with local data and compatibility constraints. The question is not how such data are generated, but how they should be evaluated: among the many possible ways in which local information might be lifted, which lifts are admissible, and how should incompatibility be measured?
From this perspective, evaluation itself may be understood as a lifting problem. A presheaf induced by the given structure and local data admits, in general, many possible lifts to globally consistent objects. Choosing an evaluation scheme amounts to specifying which lifts are regarded as compatible, and how failures of compatibility are detected. In categorical terms, this corresponds to selecting a target object together with a notion of admissible morphisms, so that consistency is assessed relative to that choice.
In the present framework, this choice is encoded by a cellular sheaf deduction model and a grounding morphism into an ambient reference. The resulting sheaf does not merely represent data, but organizes a family of admissible evaluations of structural consistency. Relative cohomology and the associated mapping cone quantify the obstruction to lifting local assignments into the prescribed evaluation scheme, while the corresponding Laplacian spectra measure the magnitude and localization of this obstruction.
Seen in this way, the inverse problem is not to recover an underlying structure, but to determine how a fixed structure fails to lift under a chosen categorification of consistency. The spectral diagnostics developed in this work may therefore be viewed as decategorified summaries of a lifting problem: they transform categorical obstructions into computable, quantitative indicators that evaluate structural consistency relative to the prescribed sheaf-theoretic model.
4.3.1 Inverse Problems as Categorification
A complementary categorical interpretation, in which this verification problem is viewed as a lifting problem for sheaves regarded as fibrations, is provided in Appendix A.6. That perspective serves as conceptual background and motivation, but is not required for the constructions or spectral analyses developed in the main body of the paper.
4.3.2 Inverse Problems as a Concrete Construction
From an inverse-problem perspective, the unknown quantity is not an underlying generative mechanism, but the degree and nature of incompatibility between observed local assignments and a fixed relational structure equipped with an evaluation scheme. Given a cell complex $K$, local data, and the induced cellular sheaf $\mathcal{F}$ with restriction maps encoding admissible local relations, the inverse problem consists of diagnosing which compatibility conditions fail, at what scale, and in what manner.
Formally, the data determine a 0-cochain in $C^{0}(K;\mathcal{F})$. Local violations of compatibility correspond to nonvanishing energy under the degree-0 sheaf Laplacian, while intrinsic global obstructions correspond to nontrivial harmonic modes of the degree-$1$ Laplacian. To evaluate consistency relative to a prescribed reference, we introduce a grounding morphism $\epsilon\colon\mathcal{F}\to\mathcal{W}$ into an ambient sheaf and study the associated mapping cone complex. A nontrivial spectral response of the cone Laplacian then certifies a relative obstruction: the local assignments are mutually compatible on $K$ but fail to admit a lift that is compatible with the chosen grounding.
In this formulation, verification of structural consistency reduces to an inverse spectral problem. Rather than reconstructing or modifying the underlying structure, we infer the presence, magnitude, and localization of obstructions by analyzing the spectra and eigenmodes of a family of Laplacians. The resulting spectral witnesses provide quantitative diagnostics that decategorify the underlying lifting problem, transforming categorical obstructions to sheafification into computable scalar indicators and localized signals.
From a categorical viewpoint, it is convenient to regard a grounded cellular sheaf as an object of the slice category ${\mathbf{CellShv}}(K)/\mathcal{W}$, that is, a morphism $\epsilon\colon\mathcal{F}\to\mathcal{W}$. Within this perspective, the base and mapping-cone Laplacians define spectral invariants of the chosen evaluation scheme, and comparisons between their spectra quantify how a fixed structure fails to lift under that scheme.
Experimental regime.
In all diagnostic experiments in Section 7, grounding is performed into a common constant sheaf concentrated in degree 0. In this regime, the cone Laplacian admits an exact block decomposition (Corollary 1), and the cone spectral channels isolate grounding-induced effects without coupling to higher-degree structure.
5 Spectral witnesses of inconsistency
We introduce a spectral filtration that quantifies counterfactual inconsistency through constraint slack, rather than through changes in the underlying topology. The construction applies uniformly to both the base sheaf Laplacian and its cone (grounded) variant, and to degrees $j\in\{0,1\}$.
5.1 Equivalence Classes of Spectral Stability
Classical sheaf cohomology provides binary indicators of consistency: the presence or absence of nontrivial cohomology groups detects exact global compatibility or obstruction. In contrast, the spectral constructions introduced in this work replace this dichotomy with a graded diagnostic structure, in which consistency is measured relative to an energy scale. Rather than classifying models as simply consistent or inconsistent, spectral gaps, low-energy eigenspaces, and their grounded (mapping-cone) counterparts quantify how far a system lies from exact consistency and where this deviation is localized.
The goal of this subsection is to formalize this relaxation by organizing spectral information into filtration-valued objects that interpolate between exact cohomology and approximate consistency, and that admit principled comparison across different Laplacians and grounding regimes.
Our approach proceeds in three steps:
  1. 1.
    introduce a spectral filtration that relaxes exact cohomology into a graded notion of approximate consistency, applicable uniformly to regular and grounded (mapping-cone) Laplacians;
  2. 2.
    define a comparability notion for such filtrations via interleavings, allowing different spectral constructions to be related without appealing to analytic perturbation theory;
  3. 3.
    justify the use of scalar spectral indicators as stable summaries of these filtrations, in the sense that they respect the induced order structure and can be meaningfully compared across models and grounding regimes.
Filtration-valued objects.
Fix a finite-dimensional inner product space $C^{j}$, and let ${\rm Sub}(C^{j})$ denote the poset of linear subspaces of $C^{j}$ ordered by inclusion. Let $(\mathbb{R}_{\geq 0},\leq)$ be the usual poset.
Remark 4 (Relation to Morse-theoretic viewpoints.).
Since our analysis is carried out on a fixed finite simplicial complex equipped with cellular sheaves, all topological variation is already discretized and organized by the underlying filtration parameters. In this regime, the role played by Morse functions and handle attachments in smooth settings is subsumed by combinatorial filtrations and barcode-type structures, and no additional Morse-theoretic machinery is required for the constructions or interpretations developed here.
Definition 3 (Admissible objects and induced filtrations).
An admissible object in degree $j$ is any datum $X$ for which there is a canonically associated order-preserving map (equivalently, a functor)
$\mathcal{H}^{j}_{\bullet}(X)\;:\;(\mathbb{R}_{\geq 0},\leq)\longrightarrow\mathrm{Sub}(C^{j}),\qquad\delta\longmapsto\mathcal{H}^{j}_{\delta}(X),$
satisfying $\delta_{1}\leq\delta_{2}\Rightarrow\mathcal{H}^{j}_{\delta_{1}}(X)\subseteq\mathcal{H}^{j}_{\delta_{2}}(X)$. We write $\mathfrak{X}_{j}$ for a chosen class of such objects, and identify $X$ with its induced filtration $\mathcal{H}^{j}_{\bullet}(X)$ whenever convenient.
Remark 5.
The term admissible is used here purely as a local convention to denote objects for which such a filtration is well-defined.
5.1.1 Spectral Filtration and Almost Non-Exactness
Let $X\in\mathfrak{X}_{j}$ be an admissible object in degree $j$ in the sense of Definition 3, with induced monotone filtration
$\mathcal{H}^{j}_{\bullet}(X):(\mathbb{R}_{\geq 0},\leq)\to\mathrm{Sub}(C^{j}).$
Definition 4 ($\delta$-harmonic space).
For any threshold $\delta\geq 0$, we define the $\delta$-harmonic space of $X$ in degree $j$ by
$\mathcal{H}^{j}_{\delta}(X)\;\in\;\mathrm{Sub}(C^{j}).$
The family $\{\mathcal{H}^{j}_{\delta}(X)\}_{\delta\geq 0}$ is referred to as the spectral filtration induced by $X$.
Spectral realization in the PSD case.
When $X$ is a symmetric positive semidefinite operator $L_{j}$ on $C^{j}$ (e.g. a cellular sheaf Laplacian or a mapping cone Laplacian), with orthogonal eigenspace decomposition $C^{j}=\bigoplus_{\lambda}E_{\lambda}$, the filtration recovers the familiar form
$\mathcal{H}^{j}_{\delta}(L_{j})\;=\;\bigoplus_{\lambda\leq\delta}E_{\lambda}.$
When $\delta=0$, the space $\mathcal{H}^{j}_{0}$ coincides with the classical harmonic space $\ker L_{j}\cong H^{j}(\mathcal{F})$. For small positive $\delta$, the space $\mathcal{H}^{j}_{\delta}$ consists of near-harmonic modes, namely cochains whose inconsistency energy is small but nonzero.
All subsequent constructions reduce to the standard spectral ones in this case.
Definition 5 (Almost non-exactness).
An admissible object $X\in\mathfrak{X}_{j}$ is said to be almost non-exact in degree $j$ if
$\mathcal{H}^{j}_{0}(X)=\{0\}\quad\text{but}\quad\mathcal{H}^{j}_{\delta}(X)\neq\{0\}\ \text{for all sufficiently small }\delta>0.$
In degree 0, this means that no nontrivial global section exists, yet the failure of global existence is spectrally unstable. In degrees $j\geq 1$, it means that no obstruction to extension exists cohomologically, yet extension fails arbitrarily weakly in the spectral sense.
5.1.2 Interleaving Compatibility of Spectral Filtrations
To compare such filtrations across different constructions, we adopt a stability notion inspired by persistence theory.
Definition 6 (Interleaving).
Two filtrations $\{\mathcal{H}^{j}_{\delta}\}$ and $\{\mathcal{H}^{\prime j}_{\delta}\}$ are said to be $\eta$-interleaved if, for all $\delta\geq 0$,
$\mathcal{H}^{j}_{\delta}\;\subseteq\;\mathcal{H}^{\prime j}_{\delta+\eta},\qquad\mathcal{H}^{\prime j}_{\delta}\;\subseteq\;\mathcal{H}^{j}_{\delta+\eta}.$
We denote the $\eta$-interleaved relationship by $\mathcal{H}^{j}_{\delta}\overset{\eta}{\approx}\mathcal{H}^{\prime j}_{\delta}$.
Interleaving captures the idea that two spectral filtrations agree up to a bounded shift in spectral resolution.
5.1.3 Stable equivalence of Spectral Indicators
Remark 6 (Scope of stability).
We do not study stability under operator perturbations (e.g. Weyl-type bounds). Instead, stability is defined internally, relative to the induced spectral filtration.
Fix a degree $j$ and a finite-dimensional inner product space $C^{j}$. Let $\mathrm{Sub}(C^{j})$ denote the poset of linear subspaces of $C^{j}$ ordered by inclusion. Let $\mathfrak{X}_{j}$ be a class of admissible objects equipped with monotone filtrations
$\mathcal{H}^{j}_{\bullet}(X):(\mathbb{R}_{\geq 0},\leq)\to\mathrm{Sub}(C^{j}),\qquad\delta\mapsto\mathcal{H}^{j}_{\delta}(X).$
Definition 7 (Spectral indicator).
A spectral indicator in degree $j$ is a map
$\Phi:\mathrm{Sub}(C^{j})\to\mathbb{R}$
that is monotone under inclusion: if $U\subseteq V$ then $\Phi(U)\leq\Phi(V)$. For each $X\in\mathfrak{X}_{j}$, the associated indicator profile is the monotone function
$\Phi_{X}:\mathbb{R}_{\geq 0}\to\mathbb{R},\qquad\Phi_{X}(\delta):=\Phi\!\big(\mathcal{H}^{j}_{\delta}(X)\big).$
Definition 8 ($\eta$-stable equivalence of spectral indicators).
Let $\Phi$ and $\Psi$ be spectral indicators in degree $j$, and let $\eta\geq 0$. We say that $\Phi$ and $\Psi$ are $\eta$-stably equivalent if for every admissible object $X\in\mathfrak{X}_{j}$ their indicator profiles are $\eta$-interleaved, i.e.
$\forall\delta\geq 0:\qquad\Phi_{X}(\delta)\leq\Psi_{X}(\delta+\eta)\quad\text{and}\quad\Psi_{X}(\delta)\leq\Phi_{X}(\delta+\eta).$
In this case we write $\Phi\overset{\eta}{\approx}\Psi$.
Remark 7.
The definition uses only monotonicity and the induced filtration $\mathcal{H}^{j}_{\bullet}(X)$.
Also, all indicators considered in this work are spectral indicators in the sense of 7, including $\dim\mathcal{H}^{j}_{\delta}$, spectral gap functions, and integrated spectral witnesses.
Lemma 2 (Interleaving compatibility of spectral indicators).
Let $\mathcal{F}$ and $\mathcal{F}^{\prime}$ be two sheaf constructions whose $\delta$-harmonic filtrations are $\eta$-interleaved in degree $j$. Then a spectral indicator function $\Phi$ satisfy
$\Phi\big(\mathcal{H}^{j}_{\delta}(L_{\mathcal{F}})\big)\;\leq\;\Phi\big(\mathcal{H}^{j}_{\delta+\eta}(L_{\mathcal{F}^{\prime}})\big),\qquad\Phi\big(\mathcal{H}^{j}_{\delta}(L_{\mathcal{F}^{\prime}})\big)\;\leq\;\Phi\big(\mathcal{H}^{j}_{\delta+\eta}(L_{\mathcal{F}})\big),$
for all $\delta\geq 0$.
Proof.
By interleaving, we have $\mathcal{H}^{j}_{\delta}(L_{\mathcal{F}})\subseteq\mathcal{H}^{j}_{\delta+\eta}(L_{\mathcal{F}^{\prime}})$. Monotonicity of $\Phi$ under inclusion yields the first inequality; the second follows symmetrically. ∎
Concrete instances
A particularly simple instance of such an indicator is
$\Phi\big(\mathcal{H}^{j}_{\delta}(L_{\mathcal{F}})\big)={\rm dim}\,\mathcal{H}^{j}_{\delta}(L_{\mathcal{F}})$
which records the number of near-harmonic modes below spectral threshold $\delta$.
This argument provides a principled notion of equivalence between spectral indicators: indicators differing only by a bounded spectral shift represent the same qualitative degree of inconsistency.
5.1.4 Compatibility with Mapping Cones
The grounded diagnostics introduced earlier arise from applying a mapping cone construction to a grounding morphism $\varepsilon:\mathcal{F}\to\mathcal{W}$.
Corollary 1 (cone grounding to the fixed constant sheaf concentrated in degree 0).
Let
$\mathcal{F}\overset{\varepsilon}{\rightarrow}\mathcal{W}\overset{\varepsilon^{\prime}}{\leftarrow}\mathcal{F}^{\prime}$
be a cospan of grounding sheaves with $\mathcal{W}$ a common constant cellular sheaf concentrated in degree 0, i.e.
$C^{0}(\mathcal{W})=W,\qquad C^{k}(\mathcal{W})=0\ \text{for all }k>0,\qquad d_{\mathcal{W}}=0,$
for some fixed vector space $W$.
Let $L^{\mathrm{cone}}_{j}(\mathcal{F},\epsilon)$ and $L^{\mathrm{cone}}_{j}(\mathcal{F}^{\prime},\epsilon^{\prime})$ denote the corresponding cone Laplacians. Then the following statements hold:
  1. 1.
    For every degree $j\geq 1$, the cone Laplacians coincide with the base Laplacians up to the standard cone degree shift: L^cone_j(F,ϵ)    L_j+1(F),   L^cone_j(F’,ϵ’)    L_j+1(F’). In particular, for $j\geq 1$ the $\delta$-harmonic filtrations of the cones depend only on $\mathcal{F}$ and $\mathcal{F}^{\prime}$, and are independent of the grounding maps.
  2. 2.
    In degree $j=0$, the cone Laplacians decompose as L^cone_0(F,ϵ) = (L1(F)+ϵTϵ00 ϵϵT),   L^cone_0(F’,ϵ’) = (L1(F’)+ϵ’Tϵ’ 00 ϵ’ϵ’T), with respect to the decomposition $C^{1}(\mathcal{F})\oplus W$.
Proof.
By assumption, the claim is immediately deduced from the cone reduction A.4. ∎
Relation to general cone reductions.
In full generality, cone Laplacians may couple base and grounding degrees in a nontrivial manner. Under strong commutativity and compatibility assumptions, this coupling disappears and the cone spectrum reduces to an additive combination of base Laplacians and grounding penalties. Appendix A.4 records an idealized structural regime in which cone Laplacians decouple exactly into base and grounding components. This result is included for completeness and is not used in the experimental analysis.
In the remainder of this section, we focus on the common-ground regime relevant to our experiments and many practical modeling scenarios, where the grounding target is a fixed constant sheaf concentrated in degree 0. In this setting, the cone construction simplifies substantially, yielding an exact block decomposition that directly explains the behavior of the cone spectral channels.
The importance of this regime is not that it introduces new homological phenomena, but that it allows one to separate intrinsic inconsistency, encoded by the absolute Laplacian of $\mathcal{F}$, from counterfactual inconsistency arising from the interaction between $\mathcal{F}$ and the grounding map. In particular, the block form of Corollary 1 makes it possible to formulate this separation directly at the level of $\delta$-harmonic filtrations.
To make this separation precise, we explicitly state as a theorem that Grounding-induced obstruction is detected exactly by those intrinsic harmonic modes that are annihilated by grounding, i.e., nonzero elements of $\ker\epsilon^{\ast}\cap\ker L_{j}(\mathcal{F})$, and that this distinction persists at the filtration level rather than merely at the level of cohomology.
Theorem 1 (Separation of intrinsic and relative obstruction for degree-0 common ground).
Let $\epsilon:\mathcal{F}\to\mathcal{W}$ be a morphism of cellular sheaves over a finite simplicial complex $K$, with $\mathcal{W}=\underline{W}$ the degree-0 concentrated constant sheaf supported only at the cone apex with stalk $W$.
Furthermore let $L_{1}(\mathcal{F})$ denote the degree-$1$ sheaf Laplacian of $\mathcal{F}$, and $L^{\mathrm{cone}}_{0}(\mathcal{F},\epsilon)$ the degree-0 cone Laplacian, and define the grounding penalty operator
$G_{\epsilon}:=\epsilon^{T}\epsilon\;:\;C^{1}(K;\mathcal{F})\to C^{1}(K;\mathcal{F}).$
Then the following statements hold.
  1. 1.
    Structural reduction. In the degree-0 common-ground regime, the counterfactual cone Laplacian is unitarily equivalent to a block-diagonal operator of the form
    $L^{\mathrm{cone}}_{0}(\mathcal{F},\epsilon)\;\simeq\;\begin{pmatrix}L_{1}(\mathcal{F})+G_{\epsilon}&0\\ 0&\epsilon\epsilon^{T}\end{pmatrix}$.
    Consequently, the $\delta$-harmonic filtration of the cone channel decomposes as
    $\mathcal{H}_{\delta}\!\left(L^{\mathrm{cone}}_{0}(\mathcal{F},\epsilon)\right)\;\cong\;\mathcal{H}_{\delta}\!\left(L_{1}(\mathcal{F})+G_{\epsilon}\right)\;\oplus\;\mathcal{H}_{\delta}\!\left(\epsilon\epsilon^{T}\right),$
    up to the cone degree shift.
  2. 2.
    Separation of obstructions. Let $H_{\mathrm{intr}}:=\ker L_{1}(\mathcal{F})$ denote the space of intrinsic harmonic modes. The following are equivalent:
    1. (a)
      $\ker L^{\mathrm{cone}}_{0}(\mathcal{F},\epsilon)\neq\{0\}$;
    2. (b)
      there exists a nonzero intrinsic harmonic mode annihilated by grounding, i.e. $\ker\!\left(\epsilon|_{H_{\mathrm{intr}}}\right)\neq\{0\}$;
    3. (c)
      $\ker\!\left(L_{1}(\mathcal{F})+G_{\epsilon}\right)\cap H_{\mathrm{intr}}\neq\{0\}$.
    In particular, grounding-induced obstruction appears if and only if grounding kills an intrinsic harmonic mode.
  3. 3.
    Filtration-level gap dichotomy. If $\epsilon|_{H_{\mathrm{intr}}}$ is injective, then there exists $\gamma>0$ such that
    $\mathcal{H}_{\delta}\!\left(L^{\mathrm{cone}}_{0}(\mathcal{F},\epsilon)\right)=\{0\}\quad\text{for all }0\leq\delta<\gamma.$
    If $\epsilon|_{H_{\mathrm{intr}}}$ is not injective, then
    $\mathcal{H}_{\delta}\!\left(L^{\mathrm{cone}}_{0}(\mathcal{F},\epsilon)\right)\neq\{0\}\quad\text{for all }\delta>0.$
  4. 4.
    Strict separation from intrinsic diagnostics. There exist pairs of grounding morphisms $\epsilon,\epsilon^{\prime}$ on the same sheaf $\mathcal{F}$ such that the intrinsic Laplacian $L_{1}(\mathcal{F})$ (and hence the intrinsic harmonic space) is identical for both, while
    $\ker L^{\mathrm{cone}}_{0}(\mathcal{F},\epsilon)=\{0\},\qquad\ker L^{\mathrm{cone}}_{0}(\mathcal{F},\epsilon^{\prime})\neq\{0\}.$
    Thus intrinsic diagnostics alone cannot, in general, detect counterfactual inconsistency.
Proof.
It directly follows from the structure reduction (i.e., Corollary 1) and the induced long exact sequence of relative sheaf cohomologies (i.e., Lemma 1. ∎
5.2 Feasible Sets and Variational Characterization
Let $K$ be a finite simplicial complex and $\mathcal{F}$ a cellular sheaf on $K$. For $j\in\{0,1\}$, let $L_{j}\;=\;d_{j-1}d_{j-1}^{\top}+d_{j}^{\top}d_{j}$ denote the degree-$j$ sheaf Laplacian acting on $C^{j}(K;\mathcal{F})$ and $L_{j}^{\mathrm{cone}}$ the corresponding Laplacian induced by the cone construction that enforces ambient grounding.
Definition 9 (Consistency energy).
For $j\in\{0,1\}$ and a cochain $x$ in the appropriate cochain space, define the consistency energy
$E_{j}(x)\;:=\;\langle x,L_{j}x\rangle,\qquad E_{j}^{\mathrm{cone}}(x)\;:=\;\langle x,L_{j}^{\mathrm{cone}}x\rangle.$
Interpretation. For $j=0$, the energy $E_{0}$ measures disagreement of node-level assignments under sheaf restrictions, while $E_{0}^{\mathrm{cone}}$ additionally measures violation of ambient grounding. For $j=1$, the energy $E_{1}$ measures residual obstruction (twists) after local consistency, and $E_{1}^{\mathrm{cone}}$ measures obstructions that persist even under grounding.
Because all Laplacians considered are symmetric positive semidefinite, the above energies induce a natural feasibility filtration.
Definition 10 ($\epsilon$-feasible set).
For $j\in\{0,1\}$ and $\epsilon\geq 0$, define
$\mathcal{S}_{j}(\epsilon)\;:=\;\big\{\,x\in C^{j}(K;\mathcal{F})\;:\;E_{j}(x)\leq\epsilon\,\big\},$
and analogously for the cone Laplacian,
$\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon)\;:=\;\big\{\,x\;:\;E_{j}^{\mathrm{cone}}(x)\leq\epsilon\,\big\}.$
Remark 8 (Feasibility filtration).
For each $j\in\{0,1\}$, the families $\{\mathcal{S}_{j}(\epsilon)\}_{\epsilon\geq 0}$ and $\{\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon)\}_{\epsilon\geq 0}$ form nested filtrations:
$0\leq\epsilon_{1}\leq\epsilon_{2}\quad\Longrightarrow\quad\mathcal{S}_{j}(\epsilon_{1})\subseteq\mathcal{S}_{j}(\epsilon_{2}),\qquad\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon_{1})\subseteq\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon_{2}).$
Interpretation. The parameter $\epsilon$ represents a budget of inconsistency. Increasing $\epsilon$ allows progressively larger violations of consistency or grounding constraints, yielding a persistence-style notion of feasibility.
Lemma 3 (Courant–Fischer Min–Max Theorem [1]).
Let $L_{j}$ have eigenvalues
$0=\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\cdots.$
Then
$\lambda_{1}=\min_{\begin{subarray}{c}x\neq 0\\ x\perp\ker L_{j}\end{subarray}}\frac{E_{j}(x)}{\|x\|^{2}},$
and $\lambda_{k}$ is the smallest value of $\epsilon$ for which $\mathcal{S}_{j}(\epsilon)$ contains a $k$-dimensional linear subspace modulo $\ker L_{j}$. An analogous statement holds for $L_{j}^{\mathrm{cone}}$.
Proof.
Direct application of Min-Max theorem [1]. ∎
Interpretation. Eigenvalues represent the critical slack levels at which new feasible directions appear. Small eigenvalues correspond to weakly violated constraints, while large eigenvalues correspond to strongly violated constraints.
5.3 Integrated and Local Spectral Witnesses
5.3.1 Global Spectral Witness
Rather than relying on a single eigenvalue, we aggregate information across the feasibility filtration.
Definition 11 (Global Spectral Witness).
Fix $0\leq\epsilon_{0}<\epsilon_{1}$. For $j\in\{0,1\}$, define (global) spectral witness at degree $j$ by
$\mathcal{I}_{j}\;:=\;\int_{\epsilon_{0}}^{\epsilon_{1}}\Big(\min_{\begin{subarray}{c}x\neq 0\\ x\perp\ker L_{j}\end{subarray}}\frac{E_{j}(x)}{\|x\|^{2}}\;\wedge\;\epsilon\Big)\,d\epsilon,$
and analogously for the cone Laplacian, (global) coned spectral witness at degree $j$ is
$\mathcal{I}_{j}^{\mathrm{cone}}\;:=\;\int_{\epsilon_{0}}^{\epsilon_{1}}\Big(\min_{\begin{subarray}{c}x\neq 0\\ x\perp\ker L_{j}^{\mathrm{cone}}\end{subarray}}\frac{E_{j}^{\mathrm{cone}}(x)}{\|x\|^{2}}\;\wedge\;\epsilon\Big)\,d\epsilon,$
where $a\wedge b:=\min\{a,b\}$.
Interpretation. The quantity $\mathcal{I}_{j}$ measures how long infeasibility persists under increasing slack at degree $j$. The cone version $\mathcal{I}_{j}^{\mathrm{cone}}$ measures persistence of infeasibility once ambient grounding is enforced. Integrating over $\epsilon$ yields a persistence-style invariant without altering the base topology.
Remark 9 (Role of the kernel and spectral gaps).
The dimension of $\ker L_{j}$ encodes the exact cohomology group $H^{j}(K;\mathcal{F})$. Accordingly, all spectral witnesses are defined on the orthogonal complement of the kernel in order to factor out gauge symmetries associated with exact consistency. As a consequence, spectral witnesses measure deviation from exact consistency rather than consistency itself. In particular, a small spectral witness should not be interpreted as evidence of consistency, but rather as indicating that any inconsistency present is weak or poorly resolved at the given spectral scale.
Summary
The base Laplacians $(L_{0},L_{1})$ quantify ungrounded consistency and obstruction, while the cone Laplacians $(L_{0}^{\mathrm{cone}},L_{1}^{\mathrm{cone}})$ quantify grounded feasibility. The $\epsilon$-feasible sets induce a spectral filtration whose critical values are eigenvalues, and the resulting spectral witnesses provide a quantitative measure of counterfactual inconsistency.
5.3.2 Local spectral witnesses
Global spectral witnesses certify the existence and persistence of inconsistency, but they do not identify its spatial origin. To localize inconsistency to specific cells, we refine the $\epsilon$-feasible spectral filtration by decomposing consistency energy into cellwise contributions associated with low-energy spectral modes.
Definition 12.
Let $L_{j}$ be the degree-$j$ (regular or cone) Laplacian with eigenpairs
$L_{j}v_{\ell}=\lambda_{\ell}v_{\ell},\qquad 0=\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\cdots.$
Fix $\epsilon>0$ and define the spectral feasible subspace
$\mathcal{V}_{j}(\epsilon)\;:=\;\mathrm{span}\{\,v_{\ell}\;:\;\lambda_{\ell}\leq\epsilon\,\}\;\subseteq\;C^{j}(K;\mathcal{F}),$
with the understanding that the kernel $\ker L_{j}$ is excluded when $j>0$.
Recall that the consistency energy admits the Hodge decomposition
$E_{j}(x)=\|d_{j}x\|^{2}+\|d_{j-1}^{\top}x\|^{2},$
where the second term is absent for $j=0$. Each term decomposes additively over the (j+1)(j{+}1)-cells and (j1)(j{-}1)-cells, respectively.
Definition 13 (Local spectral witness).
Let $L_{j}$ be the degree-$j$ sheaf Laplacian on $C^{j}(K;\mathcal{F})$, with eigendecomposition
$L_{j}v_{\ell}=\lambda_{\ell}v_{\ell},\qquad\lambda_{\ell}\geq 0,$
and let $w:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ be a nonnegative weighting function supported on a low-energy regime. For a spectral threshold $\epsilon>0$, we define the local spectral witness
$W_{j,\epsilon}:K_{j}\longrightarrow\mathbb{R}_{\geq 0}$
by
$W_{j,\epsilon}(e):=\sum_{\lambda_{\ell}\leq\epsilon}w(\lambda_{\ell})\left(\sum_{\begin{subarray}{c}c_{j+1}\in K_{j+1}\\ c_{j+1}\geq e\end{subarray}}\bigl\|(d_{j}v_{\ell})[c_{j+1}]\bigr\|^{2}\;+\;\sum_{\begin{subarray}{c}c_{j-1}\in K_{j-1}\\ c_{j-1}\leq e\end{subarray}}\bigl\|(d_{j-1}^{T}v_{\ell})[c_{j-1}]\bigr\|^{2}\right).$
Here $(d_{j}v_{\ell})[c_{j+1}]\in\mathcal{F}(c_{j+1})$ and $(d_{j-1}^{T}v_{\ell})[c_{j-1}]\in\mathcal{F}(c_{j-1})$ denote the components of $d_{j}v_{\ell}\in C^{j+1}(K;\mathcal{F})$ and $d_{j-1}^{T}v_{\ell}\in C^{j-1}(K;\mathcal{F})$, respectively, under the canonical direct-sum decompositions of the cochain spaces.
An analogous definition applies to the cone Laplacian $L_{j}^{\mathrm{cone}}$, yielding $\mathcal{W}_{j,\epsilon}^{\mathrm{cone}}(c)$.
Remark 10.
The local spectral witness aggregates contributions from multiple low-energy eigenmodes, making it more stable than localization based on a single eigenvector. This mirrors the use of persistence in topological data analysis, but operates entirely within a fixed cellular complex.
Interpretation
  • For $j=0$, $\mathcal{W}_{0,\epsilon}(c)$ localizes inconsistency to edges (and, by aggregation, to nodes), measuring which local constraints are violated by low-energy modes.
  • For $j=1$, $\mathcal{W}_{1,\epsilon}(c)$ localizes obstruction both to (j+1)(j{+}1)-cells (triangle-level twists via $d_{1}$) and to edges (divergence or gauge components via $d_{0}^{\top}$).
  • Cone versions $\mathcal{W}_{j,\epsilon}^{\mathrm{cone}}$ localize inconsistency that persists under ambient grounding, isolating defects that are invisible to ungrounded analysis.
As $\epsilon$ increases, additional spectral modes contribute, yielding a persistence-style evolution of localized inconsistency. In this sense, $\mathcal{W}_{j,\epsilon}$ provides a cell-level refinement of the global spectral witness, identifying not only whether inconsistency exists, but where it is concentrated.
6 Taxonomy of Inconsistency Mechanisms
Table 1 shows that distinct inconsistency mechanisms activate different spectral channels, and that counterfactual inconsistency is uniquely detected by the degree-0 mapping-cone Laplacian, which computes relative rather than intrinsic cohomology. Each inconsistency mechanism activates at most one dominant spectral channel.
 [^1] We supplementary records the ground penalty $P_{\rm gr}=\epsilon\epsilon^{T}$ for a grounding stress that expresses the failure of local sections to embed compatibly into the ambient reference sheaf at dimension 0. This measures tension introduced by grounding without necessarily inducing a topological obstruction.
Table 1: Taxonomy of inconsistency mechanisms and their spectral diagnostics Each inconsistency mechanism is associated with a specific cochain degree $j$ and the corresponding operator. The corresponding spectral channel detects failures of consistency at different structural levels.
Channel Operator Diagnostic meaning
Local feasibility $L_{0}(\mathcal{F})$ Local inconsistency / constraint violation
Intrinsic obstruction $L_{1}(\mathcal{F})$ Global intrinsic inconsistency (absolute obstruction)
Grounding-induced obstruction $L^{\mathrm{cone}}_{0}\;\equiv\;L_{1}(\mathcal{F})+\epsilon^{T}\epsilon$ Relative inconsistency induced by grounding
Ground utilization penalty $\epsilon\epsilon^{T}$ Unused or weakly engaged grounding directions (auxiliary diagnostic)
Note. In the degree-0 common-ground regime, the cone Laplacian splits into an intrinsic-augmented block $L_{1}(\mathcal{F})+\epsilon^{T}\epsilon$ and a ground-only penalty block $\epsilon\epsilon^{T}$. Only the former corresponds to a cone-degree Hodge Laplacian; the latter is reported as an auxiliary grounding diagnostic.
7 Diagnostic Experiments
All experiments are comparative: each consists of a base and a controlled construction that are indistinguishable under weaker diagnostics but separable under the proposed sheaf-theoretic spectral criteria. No learning is performed; all quantities are computed directly from cellular sheaf Laplacians. We organize experiments according to the discriminative characteristic they demonstrate: existence, magnitude, localization, and relativity, as in the Table 1.
In the degree-0 common-ground regime used throughout the experiments, the mapping cone Laplacian admits a block decomposition into an intrinsic-augmented operator $L_{1}(\mathcal{F})+\epsilon^{T}\epsilon$ and a ground-only penalty operator $\epsilon\epsilon^{T}$. We refer to the former as the relative cone channel, as it detects relative inconsistency arising from the interaction between intrinsic structure and grounding. The latter is not a sheaf Laplacian, but a positive semidefinite Gram operator measuring the utilization of the ground space, and is reported as an auxiliary diagnostic.
Accordingly, experimental results report three Laplacian-based spectral channels ($L_{0}(\mathcal{F})$, $L_{1}(\mathcal{F})$, $L_{0}^{\rm cone}(\mathcal{F})$) together with an additional ground utilization spectrum. This separation ensures that all Laplacian channels retain a clear Hodge-theoretic interpretation.
All results reported below are consistent with the detector summary in Table 5.
Remark on numerical scaling.
In several experiments, the raw eigenvalue scales of the (base or cone) Laplacians can vary substantially. Since our spectral witnesses integrate a clipped Rayleigh minimum over a fixed interval in the slack parameter, an uncalibrated choice of endpoints may lead to saturation and reduce discriminability. For reporting and cross-experiment comparability, we therefore apply a simple per-Laplacian rescaling (described in Appendix B) that preserves kernels and eigenvector structure and only normalizes the eigenvalue scale.
7.1 Existence: Global Consistency as Kernel Presence
We first test whether a globally consistent assignment exists. Table 2 compares a trivial line bundle and a Möbius bundle defined on the same cycle complex. The trivial bundle admits a global section, reflected by a one-dimensional kernel of the 0-th sheaf Laplacian $L_{0}$. The Möbius bundle admits no global section, and $\ker L_{0}$ is trivial. Kernel presence here constitutes an exact witness of logical consistency.
7.2 Magnitude: Degree of Inconsistency via Spectral Gap
When no global section exists up to spectral stability, we next ask how severe the inconsistency is. Table 3 compares a hidden twist and a noisy trivial bundle. Both have trivial kernel, but differ in their smallest nonzero eigenvalue and in the integrated spectral witness $\mathcal{I}_{0}$. These quantities measure the magnitude of inconsistency, i.e. the distance from admitting a global section, independent of spatial localization.
7.3 Localization: Spatial Structure of Logical Defects
Magnitude alone does not distinguish structured defects from diffuse noise. Figure 1 revisits the hidden twist and noisy trivial constructions through their low-energy eigenvectors. Despite similar global spectra, the hidden twist exhibits strongly localized eigenmodes, while the noisy trivial bundle shows diffuse energy. Localization of low-energy modes thus provides a local spectral witness of structured logical defects. Heatmaps are shown only in this setting, where the kernel vanishes and localization is the discriminative feature.
7.4 Relativity: Grounding-Induced Inconsistency via Mapping Cones
Finally, we examine inconsistency relative to a grounding or counterfactual reference. Table 4 considers two grounding morphisms applied to the same intrinsic sheaf. The regular sheaf Laplacians $L_{0}$ and $L_{1}$ are identical and have trivial kernel in both cases. However, the cone Laplacian $L^{\mathrm{cone}}_{0}$ acquires a nontrivial kernel only under rank-deficient grounding. This kernel detects a relative obstruction that is invisible to absolute sheaf cohomology, demonstrating counterfactual inconsistency.
Table 2: Existence of global consistency
Construction $\lambda_{\min}\!\left(L_{0}\right)$ $\dim\ker L_{0}$
Trivial bundle $0.0000$ 11
Möbius bundle $0.0979$ 0
Table 3: Magnitude of inconsistency
Construction $\lambda^{+}_{\min}\!\left(L_{0}\right)$ $\mathcal{I}_{0}$
Hidden twist $0.021$ $0.014$
Noisy trivial $0.086$ $0.062$
(a) Base $j=0$
(b) Base $j=1$
(c) Cone $j=0$
(d) Base $j=0$
(e) Base $j=1$
(f) Cone $j=0$
Figure 1: Local spectral witnesses for hidden twist (top row) and noisy trivial assignment (bottom row). Each row visualizes local spectral witnesses $\mathcal{W}_{j,\epsilon}$ on the same cycle complex, comparing intrinsic (base) Laplacian channels with grounded (mapping-cone) diagnostics. Shown are heatmaps induced by low-energy eigenmodes of the intrinsic $j=0$ and $j=1$ Laplacians, together with the relative cone channel corresponding to the intrinsic $j=1$ structure. Although the hidden twist and noisy trivial constructions exhibit comparable global spectral gaps, the hidden twist produces sharply localized low-energy modes, while the noisy trivial case displays diffuse energy. This localization behavior distinguishes structured logical defects from unstructured noise at the level of local spectral witnesses.
Table 4: Relative inconsistency detected by the mapping cone.
Grounding $\lambda_{\min}\!\left(L^{\mathrm{cone}}_{0}\right)$ $\dim\ker L^{\mathrm{cone}}_{0}$
Full-rank grounding $0.112$ 0
Rank-deficient grounding $0.0000$ 11
Table 5: Summary of significant spectral detectors across experiments. Each experiment is designed to isolate a single discriminative characteristic of logical inconsistency. Only detectors that are theoretically meaningful in the corresponding regime are reported.
Discriminative characteristic Experiment Primary spectral detector Interpretation
Existence Trivial vs. Möbius $\dim\ker L_{0}$ Exact existence of a global section
Magnitude Hidden Twist vs. Noisy Trivial $\lambda_{\min}^{+}(L_{0}),\ \mathcal{I}_{0}$ Degree of inconsistency (distance from exact consistency)
Localization Hidden Twist vs. Noisy Trivial Support of low-energy eigenvectors Spatial concentration of structured defects
Relativity Cone grounding comparison $\dim\ker L^{\mathrm{cone}}_{0}$ Inconsistency relative to grounding
Table 5 demonstrates that different inconsistency mechanisms activate distinct spectral channels, with relative inconsistency uniquely detected by the degree-0 mapping-cone Laplacian, corresponding to nontrivial relative cohomology.
8 Discussion and Limitations
Limitations.
The framework presented here is intentionally evaluative rather than generative. We assume that a hypothetical graph structure and a grounding map are specified a priori, and we do not address the problems of causal discovery, model learning, or intervention design. The analysis is restricted to cellular sheaves constructed on clique complexes truncated at dimension two; while this is sufficient to capture the inconsistency mechanisms studied here, higher-dimensional constructions may be required to model more complex interactions. In addition, the choice of grounding and of tolerance parameters in the construction of stalk intersections influences the resulting spectral witnesses and should be interpreted as part of the modeling assumptions rather than as universal quantities. Finally, the present work focuses on structural diagnostics rather than scalability, and no claims are made regarding computational efficiency on large or densely connected graphs.
Conclusion
We have presented a sheaf-theoretic and spectral framework for evaluating structural inconsistency in fixed relational systems under grounding. Adopting an explicitly evaluative perspective, we treat feasibility not as a generative or interventional task, but as an inverse problem: given a model specified by local relations and an ambient reference encoding admissible global behavior, determine whether they are structurally compatible, and if not, characterize the nature and localization of the obstruction.
The central conceptual contribution of this work is the formulation of structural inconsistency as a relative obstruction. By introducing a grounding morphism from a cellular sheaf of model behavior into an ambient reference sheaf and analyzing the associated mapping cone, we obtain a relative cohomology theory that separates intrinsic obstructions of the model from incompatibilities induced by grounding. This relative viewpoint refines existing sheaf-theoretic approaches to consistency, which are sensitive only to absolute cohomological obstructions, and provides a principled algebraic language for reasoning about feasibility in grounded systems.
To render these obstructions computable and diagnostically meaningful, we developed spectral witnesses derived from both regular and mapping-cone sheaf Laplacians. Rather than relying solely on kernel dimensions or exact cohomology, we used spectral gaps, integrated variational quantities, and low-energy eigenmode localization to quantify the magnitude, robustness, and spatial structure of inconsistency. These constructions yield a family of complementary spectral channels that distinguish local constraint violation, intrinsic global obstruction, and grounding-induced incompatibility within a unified Hodge-theoretic framework.
Our comparative experiments confirm that these spectral diagnostics behave in accordance with the underlying theory. Constructions that are indistinguishable under absolute cohomology or ungrounded Laplacians can be separated by mapping-cone Laplacians when relative obstructions are present. In particular, grounding-induced inconsistency arises precisely when intrinsic harmonic modes fail to embed compatibly into the ambient reference, a phenomenon that is invisible to ungrounded diagnostics.
Taken together, this work establishes mapping-cone Laplacian spectra as a principled evaluation layer for structural feasibility in sheaf-theoretic models. By decoupling diagnosis from learning, optimization, or model construction, the proposed framework complements existing approaches in sheaf-based data fusion, abstraction, and consistency analysis, and provides a mathematically grounded basis for analyzing relative feasibility in fixed systems. The extensions outlined in the following section indicate several directions in which this evaluative viewpoint may be further generalized, both theoretically and in application-driven settings.
9 Future work
Our stability notion in Section 5 is formulated at the level of filtration-valued admissible objects $X\mapsto\mathcal{H}^{j}_{\bullet}(X)$. Several natural extensions remain.
Theoretical directions.
(Spectral invariant). From a theoretical standpoint, the central open problem is to understand which spectral diagnostics are invariant — up to controlled interleaving — under changes of grounding, degree, and admissible energy. We do not aim to recover or classify relative cohomology groups beyond what is detected by the induced filtrations.
(Interleaving of global and local witnesses). In this work we defined global and local spectral witnesses derived from low-energy structure of $\mathcal{H}^{j}_{\bullet}(X)$. A first direction is to formalize global and local witnesses as interleaved spectral indicators: namely, to identify conditions under which the induced indicator profiles $\delta\mapsto\Phi_{X}(\delta)$ and $\delta\mapsto\Psi_{X}(\delta)$ associated to global and local constructions are $\eta$-interleaved (Definition 8) for a controlled shift $\eta$. Such a result would provide a principled comparability statement between global inconsistency magnitude and its localization.
(Base vs. cone channels under interleavings). A central motivation of the mapping-cone construction is that regular and coned Laplacians probe complementary obstruction types. An important next step is to characterize when the filtrations induced by regular and coned constructions are interleaved, and consequently when their global/local witnesses are interleaved spectral indicators. Concretely, this amounts to relating the base filtration $\mathcal{H}^{j}_{\bullet}(L_{j})$ and the cone filtration $\mathcal{H}^{j}_{\bullet}(L^{\mathrm{cone}}_{j})$ through quantitative hypotheses on grounding/embedding defects, thereby enabling stable comparison of intrinsic and relative obstruction channels.
(Beyond Laplacians: admissible non-linear objects). While the present paper focuses on Laplacian-based admissible objects, the stability framework only requires a canonical monotone filtration $\mathcal{H}^{j}_{\bullet}(X)$. It is therefore natural to extend the theory to non-Laplacian constructions, such as nonlinear inconsistency energies, robust penalties, or variational objectives, provided they induce well-defined low-energy subspace filtrations. A systematic development would clarify which structural properties (e.g. convexity, homogeneity, or quadratic tangents) guarantee admissibility and permit direct comparison with Laplacian-based channels.
(Applied admissible classes). Finally, it would be valuable to identify and study application-driven admissible classes in which the filtration viewpoint yields new diagnostics. Promising candidates include energy-based constraint systems arising in counterfactual evaluation, program- or logic-consistency checking on graphs, perspective-dependent grounding in multi-agent systems, and robust data-fusion pipelines. Establishing canonical filtrations in these settings would broaden the practical scope of spectral inconsistency witnesses while preserving the principled comparability guarantees afforded by interleavings.
Applied and diagnostic directions.
From an applied perspective, varying the grounding morphism $\epsilon$ induces a family of counterfactual diagnostics that can be used to stress-test assumptions about reference models or ambient constraints. Interleaving guarantees ensure that conclusions drawn from spectral witnesses are robust to moderate modeling choices, such as tolerance thresholds, approximate restrictions, or alternative groundings.
For instance, in constraint satisfaction problems on graphs, admissible objects arise from quadratic penalty energies whose low-energy subspaces encode approximately feasible assignments.
Appendix A Supplementary Result and Concept
A.1 Translation functor and sign conventions
This subsection recalls the standard translation (or shift) functor on cochain complexes and fixes the associated sign conventions used throughout the paper. All constructions in this subsection are classical and included solely to make the mapping cone computations in Section 4 precise and self-contained. Our conventions follow those in standard references such as Weibel [14] and Kashiwara–Schapira [11].
The Translation Functor.
Let $(C^{\ast},d_{C})$ be a cochain complex in an abelian category (or, concretely, a cochain complex of finite-dimensional vector spaces). For an integer $k\in\mathbb{Z}$, the translation functor
$k]:\mathrm{Ch}\to\mathrm{Ch}$
is defined by
$C[k]^{n}:=C^{n-k},\qquad d_{C[k]}^{n}:=(-1)^{k},d_{C}^{n-k}.$
We adopt this convention throughout. It agrees with the standard conventions used in triangulated categories and derived functor formalisms (see [14, 11]).
Sign convention.
To my knowledge, this sign convention originates from the very nature of the mapping cone in the triangulated category, where the translation functor encodes the compatibilities involved with $d_{\mathcal{F}},d_{\mathcal{W}}$ and $d_{\mathrm{cone}}$.
A.2 Geometric cone orientation and the independence
Orientation convention for the geometric cone.
Fix an ordering of the vertices of $K$ (e.g. the natural order in the clique complex construction), and orient each simplex $v_{0},\dots,v_{n}]$ in $K$ by that ordered vertex list. We orient the cone $\widehat{K}=\mathrm{Cone}(K)$ by declaring that the apex vertex $\ast$ is greater than every vertex of $K$, and orient each cone simplex by
$\ast,v_{0},\dots,v_{n}],$
where $v_{0},\dots,v_{n}]$ is the oriented base simplex. With this choice, the face map that forgets the apex,
$\partial_{0}[\ast,v_{0},\dots,v_{n}]=[v_{0},\dots,v_{n}],$
appears in the boundary operator with sign $(-1)^{0}=1$, while for $i\geq 1$ we have $\partial_{i}[\ast,v_{0},\dots,v_{n}]=[\ast,v_{0},\dots,\widehat{v_{i-1}},\dots,v_{n}]$ with sign $(-1)^{i}$. Equivalently, in the cellular cochain differential on $\widehat{K}$, the contribution from the cone incidence $\sigma\prec\ast\sigma$ carries a fixed incidence sign determined by the above orientation.
Independence of orientation.
Any other choice of orientation on the cone cells differs from the above by multiplying each cone cell $\ast\sigma$ by a sign $\pm 1$. This induces a diagonal change-of-basis on the $W$-summand of $C^{n}(\widehat{K};\widehat{\mathcal{F}})\cong C^{n}(K;\mathcal{F})\oplus C^{n-1}(K;\mathcal{W})$, and hence yields an isomorphic cochain complex. Therefore, the identification of the geometric cone complex with the translated algebraic mapping cone is canonical up to (cochain) isomorphism and does not depend on a particular orientation convention.
A.3 Long exact sequence of a mapping cone
The proof of Lemma 1.
The argument follows [14] verbatim; we reproduce it here for the reader’s convenience. First we see that there is a short exact sequence of cochain complexes of the form
$0\to C(K;\mathcal{W})\xrightarrow{i}C(K;\mathrm{cone}(\epsilon))\xrightarrow{q}C(K;\mathcal{F})[-1]\to 0,$
where $i(c)=(0,c)$, q(b,c)=bq(b,c)=-b and $C(K;\mathcal{F})[-1]^{n}=C(K;\mathcal{F})^{n+1}$ denotes the translated complex as before. We stick to the definition of cone differential as:
$d_{\rm cone}^{n}:=\begin{pmatrix}-d_{\mathcal{F}}&0\\ -\epsilon&d_{\mathcal{W}}\end{pmatrix}$.
We immediately observe that $i$ is injective, $q$ is surjective, $\mathrm{Im}\,i\cong\ker q$ by definition and $i$ is a cochain map since $d_{\rm cone}(0,c)=(0,d_{\mathcal{W}}c)$.
To see $q$ being a cochain map, we use the standard translation convention for cochain complexes A.1, in which the shifted complex $C(K;\mathcal{F})[-1]$ carries the differential $d_{\mathcal{F}}$. With this convention, we have $(d_{\mathcal{F}[-1]}\circ q)(b,c)=(q\circ d_{\rm cone})(b,c)=d_{\mathcal{F}}b$.
Via the diagram chasing on the diagram below, the connecting morphism appears to coincide with the induced map $(\epsilon^{\ast})_{\ast}$ on cohomology,
[…]
For a cocyle $b\in C^{n}(K;\mathcal{F})$, a lifting $(-b,c)\in q^{-1}(b)$ differentiates to $(0,\epsilon^{\ast}b+d_{\mathcal{W}}c)$, which corresponds to a representing element $\epsilon^{\ast}b$ in $H^{n}(K;\mathcal{W})$. ∎
A.4 A Cone Reduction
Lemma 4 (General cone reduction under Commuting Assumptions).
Assume there are two grounded sheaves $(\mathcal{F},\varepsilon:\mathcal{F}\to\mathcal{W})$ and $(\mathcal{F}^{\prime},\varepsilon^{\prime}:\mathcal{F}^{\prime}\to\mathcal{W}^{\prime})$ that induce a pair of $\eta$-interleaved $\delta$-harmonic filtrations, meaning that:
$\mathcal{H}^{j}_{\delta}(\mathcal{F})\overset{\eta}{\approx}\mathcal{H}^{j}_{\delta}(\mathcal{F}^{\prime})\text{ and }\mathcal{H}^{j}_{\delta}(\mathcal{W})\overset{\eta}{\approx}\mathcal{H}^{j}_{\delta}(\mathcal{W}^{\prime}).$
If they additionally suffice the following conditions:
  1. 1.
    $\varepsilon,\varepsilon^{\prime}$ intertwines $d^{T}$, namely d_W^Tε=εd_F^T, d_W’^Tε’=ε’ d_F’^T
  2. 2.
    Sheaf Laplacians and Gramian matrices of grounding maps are commutative, precisely: [L_F,ε^Tε]=[L_W,εε^T] =[L_F’,ε’^Tε’]=[L_W’,ε’ε’^T] =0,
then the corresponding filtrations obtained from their mapping cone Laplacians are $(\eta+\mathbf{v})$-interleaved, where $\mathbf{v}$ is defined by
$\mathbf{v}:=\mathbf{v}(\varepsilon,\varepsilon^{\prime})=\max_{\begin{subarray}{c}b_{i}\in\mathrm{Spec}(\varepsilon^{T}\varepsilon)\\ b^{\prime}_{j}\in\in\mathrm{Spec}(\varepsilon\varepsilon^{T})\end{subarray}}|b_{i}-b^{\prime}_{j}|.$
Furthermore, if the Gramian matrices of grounding maps are $\theta$-interleaved, meaning that
$\varepsilon^{T}\varepsilon\overset{\theta}{\approx}\varepsilon^{\prime T}\varepsilon^{\prime}\text{ and}$ (A.4)
$\varepsilon\varepsilon^{T}\overset{\theta}{\approx}\varepsilon^{\prime}\varepsilon^{\prime T},$
then we can take $\mathbf{v}=\theta$.
Implication.
Under these assumptions, the cone Laplacians admit a simultaneous spectral decomposition in which grounding contributes additively to the base operators. Consequently, interleavings between the $\delta$-harmonic filtrations of $L^{\mathrm{cone}}_{j}(\mathcal{F},\epsilon)$ and $L^{\mathrm{cone}}_{j}(\mathcal{F}^{\prime},\epsilon^{\prime})$ reduce to bounded shifts determined by the corresponding grounding penalty spectra (c.f., Corollary 1).
Proof.
The argument is a direct consequence of simultaneous diagonalization under the stated commutativity assumptions.
By calculation, we can get the decomposition of the cone Laplacian for $(\mathcal{F},\varepsilon)$:
$L^{cone}_{\mathcal{F}}=\begin{pmatrix}L_{\mathcal{F}}+\varepsilon^{T}\varepsilon&\varepsilon^{T}d_{\mathcal{W}}-d_{\mathcal{F}}\varepsilon^{T}\\ d_{\mathcal{W}}^{T}\varepsilon-\varepsilon d_{\mathcal{F}}^{T}&L_{\mathcal{W}}+\varepsilon\varepsilon^{T}\end{pmatrix}=\begin{pmatrix}L_{\mathcal{F}}+\varepsilon^{T}\varepsilon&0\\ 0&L_{\mathcal{W}}+\varepsilon\varepsilon^{T}\end{pmatrix}+\begin{pmatrix}0&\varepsilon^{T}d_{\mathcal{W}}-d_{\mathcal{F}}\varepsilon^{T}\\ d_{\mathcal{W}}^{T}\varepsilon-\varepsilon d_{\mathcal{F}}^{T}&0\end{pmatrix}$,
where the 2nd term vanishes under the condition 1. Similarly for $(\mathcal{F}^{\prime},\varepsilon^{\prime})$, the cone Laplacians have decomposition of the same form.
Now that the claim is reduced to show
  • $\mathcal{H}^{j}_{\delta}(L_{\mathcal{F}}+\varepsilon^{T}\varepsilon)\subseteq\mathcal{H}^{j}_{\delta+\eta+\mathbf{v}}(L_{\mathcal{F}^{\prime}}+\varepsilon^{\prime T}\varepsilon^{\prime})$ and $\mathcal{H}^{j}_{\delta}(L_{\mathcal{W}}+\varepsilon\varepsilon^{T})\subseteq\mathcal{H}^{j}_{\delta+\eta+\mathbf{v}}(L_{\mathcal{W}^{\prime}}+\varepsilon^{\prime}\varepsilon^{\prime T})$ and,
  • $\mathcal{H}^{j}_{\delta}(L_{\mathcal{F}^{\prime}}+\varepsilon^{\prime T}\varepsilon^{\prime})\subseteq\mathcal{H}^{j}_{\delta+\eta+\mathbf{v}}(L_{\mathcal{F}}+\varepsilon^{T}\varepsilon)$ and $\mathcal{H}^{j}_{\delta}(L_{\mathcal{W}^{\prime}}+\varepsilon^{\prime}\varepsilon^{\prime T})\subseteq\mathcal{H}^{j}_{\delta+\eta+\mathbf{v}}(L_{\mathcal{W}}+\varepsilon\varepsilon^{T})$,
since then
$\mathcal{H}^{j}_{\delta}(L^{cone}_{\mathcal{F}})$ $=\mathcal{H}^{j}_{\delta}(L_{\mathcal{F}}+\varepsilon^{T}\varepsilon)\oplus\mathcal{H}^{j}_{\delta}(L_{\mathcal{W}}+\varepsilon\varepsilon^{T})$
$\subseteq\mathcal{H}^{j}_{\delta+\eta+\mathbf{v}}(L_{\mathcal{F}^{\prime}}+\varepsilon^{\prime T}\varepsilon^{\prime})\oplus\mathcal{H}^{j}_{\delta+\eta+\mathbf{v}}(L_{\mathcal{W}^{\prime}}+\varepsilon^{\prime}\varepsilon^{\prime T})$
$=\mathcal{H}^{j}_{\delta+\eta+\mathbf{v}}(L^{cone}_{\mathcal{F}^{\prime}}).$
The reverse direction is similar.
To show the first subset relation, we note that all the operators represented as matrices here are positive semi-definite, so diagonalizable, and the condition 2 asserts that each pair of summands within $\mathcal{H}^{j}_{(-)}$ admits a basis of common eigenvectors, meaning that we have the representation of the form:
$L_{\mathcal{F}}+\varepsilon^{T}\varepsilon=P^{-1}\begin{pmatrix}a_{1}+b_{1}&0&\dots&0\\ 0&a_{2}+b_{2}&\dots&0\\ \vdots&&\ddots&\vdots&\\ 0&0&\dots&a_{r}+b_{r}\end{pmatrix}P,$
with some chosen orthogonal matrix $P$. By definition of $\mathcal{H}^{j}_{(-)}$, there is no loss of generality if we assume that the diagonal factors are ordered smaller to larger, to the bottom.
For arbitrary $\delta>0$, choose $\delta_{0}$ such that $a_{k}\leq\delta_{0}\leq\delta-b_{k}$ for some $k\leq r$, then we can find $k^{\prime}\geq k$ such that $a^{\prime}_{k^{\prime}}\leq\delta_{0}+\eta$.
With the same $k^{\prime}$, we can deduce that
$a^{\prime}_{k^{\prime}}+b^{\prime}_{k^{\prime}}$ $\leq\delta_{0}+\eta+b^{\prime}_{k^{\prime}}$
$=\delta+\eta+b^{\prime}_{k^{\prime}}-b_{k}$
$\leq\delta+\eta+\mathbf{v}.$
The exact same argument is applied for $L_{\mathcal{W}}+\varepsilon\varepsilon^{T}$, and by the construction, the converse direction is deduced in the same way. Therefore we showed
$\mathcal{H}^{j}_{\delta}(L^{cone}_{\mathcal{F}})\overset{\eta+\mathbf{v}}{\approx}\mathcal{H}^{j}_{\delta}(L^{cone}_{\mathcal{F}^{\prime}}).$
If the condition A.4 is satisfied, then in the same setting above, we have $a_{k}\leq\delta_{0}$ and $b_{k}\leq\delta-\delta_{0}$ so that we can find $k^{\prime}\geq k$ such that
$a^{\prime}_{k^{\prime}}\leq\delta_{0}+\eta,\quad b^{\prime}_{k^{\prime}}\leq\delta-\delta_{0}+\theta$
holds, showing that $\mathcal{H}^{j}_{\delta}(L^{cone}_{\mathcal{F}})\overset{\eta+\theta}{\approx}\mathcal{H}^{j}_{\delta}(L^{cone}_{\mathcal{F}^{\prime}})$.
A.5 Clique Complex
The definition of the clique (or flag) complex associated to a graph is provided in this subsection for the conventions used throughout the paper.
Definition 14 (Clique Complex).
Let $G=(V,E)$ be a finite, undirected, simple graph. A clique in $G$ is a finite subset
$\sigma=\{v_{0},\dots,v_{k}\}\subset V$
such that every pair of distinct vertices in $\sigma$ is joined by an edge in $E$.
The clique complex (or flag complex) of $G$, denoted $\mathrm{Cl}(G)$, is the abstract simplicial complex defined by
$\mathrm{Cl}(G)_{k}=\bigl\{\sigma\subset V\;\big|\;|\sigma|=k+1\text{ and }\sigma\text{ is a clique in }G\bigr\}.$
By construction, the 11–skeleton of $\mathrm{Cl}(G)$ is canonically identified with the original graph $G$. The clique complex is the maximal simplicial complex whose 11–skeleton is $G$.
Functoriality and Minimality
The assignment
$G\longmapsto\mathrm{Cl}(G)$
is functorial with respect to graph homomorphisms that preserve adjacency. In particular, any automorphism of $G$ induces a simplicial automorphism of its clique complex.
The clique complex is characterized by the following minimality property: among all simplicial complexes whose 11–skeleton is $G$, $\mathrm{Cl}(G)$ is the unique maximal one. Equivalently, a simplicial complex is a clique complex if and only if it is flag, meaning that every finite set of pairwise adjacent vertices spans a simplex.
A.5.1 Dimensional Truncation
In general, the clique complex of a graph may have arbitrarily high dimension. In this work, we restrict attention to the 22–skeleton
$\mathrm{Cl}(G)^{(2)},$
consisting of vertices, edges, and triangles.
This truncation is sufficient for the constructions considered here, since
  • 0–cochains encode vertex-level data,
  • 11–cochains encode edge-level compatibility,
  • 22–cochains encode triangular (cycle-level) consistency.
All cellular sheaves, cochain complexes, and Laplacians in the main text are defined on this truncated complex. The restriction to the 22–skeleton is made for practical simplicity and experimental clarity, rather than as a theoretical limitation. While the constructions considered here probe cohomology in degrees 0 and $1$, the underlying framework naturally extends to higher-dimensional simplicial or cellular complexes, where higher-degree cochains and obstructions are expected to play a substantive role.
A.5.2 Orientation Conventions
To define cellular coboundary operators, we fix a total ordering of the vertex set $V$. Each simplex
$v_{0},\dots,v_{k}],\qquad v_{0}<\cdots<v_{k},$
is oriented by the induced ordered basis.
With this convention, the simplicial coboundary operator is given by the standard alternating sum over faces. Any other choice of vertex ordering yields a canonically isomorphic cochain complex. Consequently, all cohomological and spectral invariants considered in this work are independent of the specific ordering chosen.
A.5.3 Remarks on Alternative Complexes
Other simplicial constructions associated to graphs, such as Vietoris–Rips or Čech complexes, are frequently used in applied topology. The clique complex is distinguished here by the fact that it depends only on adjacency data and introduces no auxiliary scale or parameter.
For the purposes of studying cellular sheaves and their Laplacians on graphs, the clique complex provides a canonical simplicial thickening of the underlying graph and will be used throughout without further comment.
A.6 A Categorical Interpretation of the Inverse Problem
At a higher level of abstraction, mathematical language admits a rigorous interpretation of a direct problem as a form of decategorification: one seeks to extract a well-understood invariant or signature from a structure-enriched object. A canonical example of decategorification replaces the category of finite-dimensional vector spaces $\mathbf{FinVect}_{k}$ by the set of natural numbers $\mathbb{N}$ via the dimension function.
In contrast, from a categorical perspective, many inverse problems may be understood as instances of categorification. In particular, lifting problems associated with fibrations admit a natural categorified interpretation. Rather than asking for the existence of a single lift, a fibration encodes a structured family of lifts organized into a fiber category, together with functorial reindexing along morphisms in the base. Solvability of a lifting problem is therefore replaced by the study of a categorical object whose objects represent lifts and whose morphisms encode coherent identifications. This viewpoint is classical in the theory of fibered categories and descent [8, 2], and underlies modern homotopy-theoretic formulations in which lifting problems are governed by higher-categorical and cohomological obstructions.
Definition 15 ([12]).
Let $p\colon E\to B$ be a functor. A morphism $g\colon e\to e^{\prime}$ in $E$ is said to be $p$-cocartesian if for every morphism $g^{\prime}\colon e\to e^{\prime\prime}$ in $E$ and every morphism $f\colon p(e^{\prime})\to p(e^{\prime\prime})$ in $B$ such that $p(g^{\prime})=f\circ p(g)$, there exists a unique morphism $h\colon e^{\prime}\to e^{\prime\prime}$ in $E$ satisfying $g^{\prime}=h\circ g$ and $p(h)=f$. The functor $p$ is called an opfibration if for every object $e\in E$ and every morphism $f\colon p(e)\to b$ in $B$, there exists a $p$-cocartesian morphism $g\colon e\to e^{\prime}$ in $E$ with $p(g)=f$. This situation is summarized by the commutative diagram
[…]
Sheaves may be canonically interpreted as (op)fibrations over the underlying site, depending on variance conventions. From this viewpoint, the task of verifying consistency in our framework is translated into a quantitative and parametric study of the classical lifting problem for sheaves regarded as fibrations. In this sense, structural inconsistency arises as a categorified obstruction to lifting, rather than as a purely set-theoretic failure.
Appendix B Implementation Details
This appendix summarizes the computational pipeline used to construct cellular sheaves, sheaf Laplacians, and spectral witnesses from attributed graphs. The implementation is intended as a concrete realization of the constructions introduced in the main text and is not optimized for large-scale deployment.
B.1 Input Data and Ambient Representation
The input consists of an undirected graph $G=(V,E)$ together with a collection of node feature matrices $\{F(v)\}_{v\in V}$. Each $F(v)\in\mathbb{R}^{D_{v}\times k_{v}}$ is interpreted as a set of local attribute vectors associated with node $v$, embedded in an ambient Euclidean space.
For each node $v$, a singular value decomposition is performed on $F(v)$. Singular vectors corresponding to non-negligible singular values are retained to form an orthonormal basis $B_{v}$. The span of $B_{v}$ defines the node stalk $\mathcal{F}(v)$. This step extracts an intrinsic local representation while discarding redundant or degenerate directions.
B.2 Clique Complex Construction
From the graph $G$, we construct its clique complex $K$, truncated to dimension two. Thus $K$ consists of vertices $K_{0}=V$, edges $K_{1}=E$, and triangles $K_{2}$ corresponding to $3$-cliques in $G$. No cells of dimension greater than two are included. This truncation reflects the fact that the obstruction mechanisms of interest arise at cochain degrees 0 and $1$.
B.3 Edge Stalks via Subspace Intersection
For each edge $e=\{u,v\}\in K_{1}$, the edge stalk $\mathcal{F}(e)$ is constructed as a geometric intersection of the node stalks $\mathcal{F}(u)$ and $\mathcal{F}(v)$. Let $B_{u}$ and $B_{v}$ denote orthonormal bases of the two node stalks. We form the matrix $M_{uv}=B_{u}^{\top}B_{v}$ and compute its singular value decomposition.
Singular vectors corresponding to singular values exceeding a prescribed tolerance define directions that are nearly aligned in both stalks. These vectors form a basis $B_{e}$ for $\mathcal{F}(e)$. The restriction maps from $\mathcal{F}(u)$ and $\mathcal{F}(v)$ to $\mathcal{F}(e)$ are implemented as coordinate projections onto this basis.
The figure 2 shows how an edge feature (i.e., a section over 1-cell) is derived as the 1-dimensional intersection of endpoints node features at the specified threshold.
  • Red Plane: The subspace spanned by $u_{1}$ and $u_{2}$ (from the $B_{u}$ matrix).
  • Blue Plane: The subspace spanned by $v_{1}$ and $v_{2}$ (from the $B_{v}$ matrix).
  • Basis Vectors: The red arrows ($u_{1},u_{2}$) lie on the red plane, and the blue arrows ($v_{1},v_{2}$) lie on the blue plane.
  • $B_{int}$ (Green Line): The thick green line represents the computed intersection vector.
Figure 2: 1D intersection $\mathcal{F}(u)\cap\mathcal{F}(v)$, each of which is spanned by two 3-dimensional vectors.
B.4 Triangle Stalks via Symmetric Soft Intersection
For each triangle $\{u,v,w\}\in K_{2}$, the triangle stalk $\mathcal{F}(uvw)$ is deduced from the three incident edge stalks. Let $B_{uv},B_{vw},B_{uw}$ denote orthonormal bases of the corresponding edge stalks. We form the orthogonal projectors
$P_{uv}=B_{uv}B_{uv}^{\top},\quad P_{vw}=B_{vw}B_{vw}^{\top},\quad P_{uw}=B_{uw}B_{uw}^{\top}.$
A symmetric alignment operator is defined by composing these projectors and symmetrizing:
$T=(P_{uv}P_{vw}P_{uw})^{\top}(P_{uv}P_{vw}P_{uw}).$
Eigenvectors of $T$ corresponding to sufficiently large eigenvalues are retained, optionally after applying a spectral exponent. Their span defines the triangle stalk $\mathcal{F}(uvw)$.
This construction yields a symmetric, parameterized notion of higher-order consistency that interpolates between permissive and rigid intersection constraints. Restriction maps from edges to triangles are again implemented as coordinate projections.
B.5 Cellular Sheaf and Coboundary Operators
The above procedures define a cellular sheaf $\mathcal{F}$ on the $2$-dimensional clique complex $K$. The cochain spaces are assembled as direct sums of stalk coordinates:
$C^{0}(K;\mathcal{F})=\bigoplus_{v\in K_{0}}\mathcal{F}(v),\quad C^{1}(K;\mathcal{F})=\bigoplus_{e\in K_{1}}\mathcal{F}(e),\quad C^{2}(K;\mathcal{F})=\bigoplus_{t\in K_{2}}\mathcal{F}(t).$
Coboundary operators $d^{0}$ and $d^{1}$ are constructed as sparse block matrices encoding the restriction maps with alternating signs determined by a fixed orientation. The degree-wise sheaf Laplacians are then given by
$L_{0}=d^{0\top}d^{0},\qquad L_{1}=d^{0}d^{0\top}+d^{1\top}d^{1}.$
B.6 Grounding and Mapping Cone Construction
To evaluate counterfactual consistency, a geometric mapping cone is constructed by adjoining a cone vertex $\ast$ to the complex and connecting it to all original vertices. The stalk at $\ast$ is chosen as an ambient reference space $W=\mathbb{R}^{d_{\max}}$, where $d_{\max}$ is the maximum ambient feature dimension across nodes.
For each node $v$, the restriction map $\rho_{v,\ast}$ embeds the node stalk basis into $W$, with zero padding if necessary. This defines a grounding map $\epsilon\colon\mathcal{F}\to\mathcal{W}$ compatible with the coboundary structure. The induced cone sheaf yields cone Laplacians whose spectra encode relative obstructions to extending sections into the ambient reference.
B.7 Spectral Witnesses
Spectral witnesses are computed from the degree-wise Laplacians using their variational characterization. Global witnesses are defined from the smallest strictly positive eigenvalues, while local witnesses are obtained by projecting low-energy eigenmodes onto cells to produce node-, edge-, or triangle-wise inconsistency scores. These quantities are used throughout the experiments to distinguish local causal defects, intrinsic topological obstructions, and counterfactual inconsistency under grounding.
Empirical Normalization
For each sheaf, we compute the degree-$j$ sheaf Laplacians $L_{j}$ and, when applicable, the corresponding cone Laplacians $L^{\mathrm{cone}}_{j}$ induced by a grounding morphism $\epsilon:\mathcal{F}\to\mathcal{W}$. Spectral witnesses are derived from the smallest strictly positive eigenvalue (the spectral gap) of these operators, restricted to the orthogonal complement of the kernel. The presence of a nontrivial kernel is interpreted as exact consistency (i.e. nonvanishing cohomology), whereas a positive spectral gap quantifies deviation from consistency.
To enable comparison across experiments with different ambient dimensions and stalk sizes, we apply a simple normalization to the spectral gaps. Specifically, each gap is rescaled by a characteristic spectral scale of the corresponding Laplacian (e.g. the trace divided by rank). This normalization ensures that witness values are comparable across settings, but it does not alter the qualitative ordering of experiments or the presence or absence of kernels. Accordingly, normalization is treated as an implementation detail rather than a substantive modeling choice; the theoretical interpretation of consistency and inconsistency rests on kernel structure and relative spectral behavior, not on absolute magnitudes.
Finally, we emphasize that the kernel dimension itself is not used as a quantitative score. Rather, it serves as a discrete indicator of exact consistency, while the spectral gap on its orthogonal complement provides a robust quantitative witness of inconsistency.

### v2

@@
 Proposition 1 (Geometric cone realizes the translated mapping cone).
-Let $K$ be a finite simplicial complex and let $\varepsilon:\mathcal{F}\to\mathcal{W}$ be a cellular sheaf morphism, by assuming the compatibility condition 4.2.1, and write $\epsilon^{\ast}:C^{\ast}(K;\mathcal{F})\to C^{\ast}(K;\mathcal{W})$ for the induced cochain map.
+Let $K$ be a finite simplicial complex and let $\epsilon:\mathcal{F}\to\mathcal{W}$ be a cellular sheaf morphism, by assuming the compatibility condition 4.2.1, and write $\epsilon^{\ast}:C^{\ast}(K;\mathcal{F})\to C^{\ast}(K;\mathcal{W})$ for the induced cochain map.
 Let $\widehat{K}$ denote the geometric cone on $K$ with apex vertex $\ast$, and let $\widehat{\mathcal{F}}$ be the cone sheaf on $\widehat{K}$ defined by \$4.2.2.
@@
 Hence by definition, we have the canonical isomorphism of vector spaces
-$\Phi:C^{n}(\widehat{K};\widehat{\mathcal{F}})\cong\Bigl(\bigoplus_{\tau\in K_{n}}\mathcal{F}(\tau)\Bigr)\ \oplus\ \Bigl(\bigoplus_{\sigma\in K_{n-1}}\mathcal{W}(\sigma)\Bigr)\cong C^{n}(K;\mathcal{F})\ \oplus\ C^{n-1}(K;\mathcal{W})=\mathrm{Cone}(\epsilon^{\ast})[1]^{n},$
-where $\Phi$ is defined by
-$\Phi^{n}(\alpha)=(\alpha|_{\mathcal{F}},\,(-1)^{n}\alpha|_{\mathcal{W}}).$
+$\Phi:C^{n}(\widehat{K};\widehat{\mathcal{F}})\cong\Bigl(\bigoplus_{\tau\in K_{n}}\mathcal{F}(\tau)\Bigr)\ \oplus\ \Bigl(\bigoplus_{\sigma\in K_{n-1}}\mathcal{W}(\sigma)\Bigr)\cong C^{n}(K;\mathcal{F})\ \oplus\ C^{n-1}(K;\mathcal{W})=\mathrm{Cone}(\epsilon^{\ast})[1]^{n}.$
 To see that $\Phi$ is a cochain isomorphism, we show that $\Phi^{n+1}\circ d_{\widehat{\mathcal{F}}}=d_{\mathrm{cone}[1]}\circ\Phi^{n}$.
+[…]
 Using the apex-first cone orientation from Appendix A.2, a direct inspection of the cellular coboundary yields
 […]
-where the alternating signs in $d_{\mathcal{F}}$ and $d_{\mathcal{W}}$ arise from incidence numbers, and the factor $(-1)^{n+1}$ arises solely from the degree shift in the identification $\Phi$.
+where the cone-cone contribution satisfies
+$(d_{\widehat{\mathcal{F}}}\alpha)(\ast\tau)$ $=\sum_{\sigma\prec\tau}[\ast\tau:\ast\sigma]\,\mathcal{W}(\sigma\prec\tau)\,\alpha_{\mathcal{W}}(\ast\sigma)$
+$=-\sum_{\sigma\prec\tau}[\tau:\sigma]\,\mathcal{W}(\sigma\prec\tau)\,y(\sigma)$
+$=-(d_{\mathcal{W}}y)(\tau),\quad\forall\tau\in K_{n}.$
 These contributions assemble to
@@
 5.1.4 Compatibility with Mapping Cones
-The grounded diagnostics introduced earlier arise from applying a mapping cone construction to a grounding morphism $\varepsilon:\mathcal{F}\to\mathcal{W}$.
+The grounded diagnostics introduced earlier arise from applying a mapping cone construction to a grounding morphism $\epsilon:\mathcal{F}\to\mathcal{W}$.
 Corollary 1 (cone grounding to the fixed constant sheaf concentrated in degree 0).
 Let
-$\mathcal{F}\overset{\varepsilon}{\rightarrow}\mathcal{W}\overset{\varepsilon^{\prime}}{\leftarrow}\mathcal{F}^{\prime}$
+$\mathcal{F}\overset{\epsilon}{\rightarrow}\mathcal{W}\overset{\epsilon^{\prime}}{\leftarrow}\mathcal{F}^{\prime}$
 be a cospan of grounding sheaves with $\mathcal{W}$ a common constant cellular sheaf concentrated in degree 0, i.e.
@@
 Proof.
-It directly follows from the structure reduction (i.e., Corollary 1) and the induced long exact sequence of relative sheaf cohomologies (i.e., Lemma 1. ∎
+It directly follows from the structure reduction (i.e., Corollary 1) and the induced long exact sequence of relative sheaf cohomologies (i.e., Lemma 1). ∎
 5.2 Feasible Sets and Variational Characterization
@@
 Because all Laplacians considered are symmetric positive semidefinite, the above energies induce a natural feasibility filtration.
-Definition 10 ($\epsilon$-feasible set).
-For $j\in\{0,1\}$ and $\epsilon\geq 0$, define
-$\mathcal{S}_{j}(\epsilon)\;:=\;\big\{\,x\in C^{j}(K;\mathcal{F})\;:\;E_{j}(x)\leq\epsilon\,\big\},$
+Definition 10 ($\delta$-feasible set).
+For $j\in\{0,1\}$ and $\delta\geq 0$, define
+$\mathcal{S}_{j}(\delta)\;:=\;\big\{\,x\in C^{j}(K;\mathcal{F})\;:\;E_{j}(x)\leq\delta\,\big\},$
 and analogously for the cone Laplacian,
-$\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon)\;:=\;\big\{\,x\;:\;E_{j}^{\mathrm{cone}}(x)\leq\epsilon\,\big\}.$
+$\mathcal{S}_{j}^{\mathrm{cone}}(\delta)\;:=\;\big\{\,x\;:\;E_{j}^{\mathrm{cone}}(x)\leq\delta\,\big\}.$
 Remark 8 (Feasibility filtration).
-For each $j\in\{0,1\}$, the families $\{\mathcal{S}_{j}(\epsilon)\}_{\epsilon\geq 0}$ and $\{\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon)\}_{\epsilon\geq 0}$ form nested filtrations:
-$0\leq\epsilon_{1}\leq\epsilon_{2}\quad\Longrightarrow\quad\mathcal{S}_{j}(\epsilon_{1})\subseteq\mathcal{S}_{j}(\epsilon_{2}),\qquad\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon_{1})\subseteq\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon_{2}).$
+For each $j\in\{0,1\}$, the families $\{\mathcal{S}_{j}(\delta)\}_{\delta\geq 0}$ and $\{\mathcal{S}_{j}^{\mathrm{cone}}(\delta)\}_{\delta\geq 0}$ form nested filtrations:
+$0\leq\delta_{1}\leq\delta_{2}\quad\Longrightarrow\quad\mathcal{S}_{j}(\delta_{1})\subseteq\mathcal{S}_{j}(\delta_{2}),\qquad\mathcal{S}_{j}^{\mathrm{cone}}(\delta_{1})\subseteq\mathcal{S}_{j}^{\mathrm{cone}}(\delta_{2}).$
 Interpretation.
-The parameter $\epsilon$ represents a budget of inconsistency.
-Increasing $\epsilon$ allows progressively larger violations of consistency or grounding constraints, yielding a persistence-style notion of feasibility.
-Lemma 3 (Courant–Fischer Min–Max Theorem [1]).
-Let $L_{j}$ have eigenvalues
-$0=\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\cdots.$
-Then
-$\lambda_{1}=\min_{\begin{subarray}{c}x\neq 0\\ x\perp\ker L_{j}\end{subarray}}\frac{E_{j}(x)}{\|x\|^{2}},$
-and $\lambda_{k}$ is the smallest value of $\epsilon$ for which $\mathcal{S}_{j}(\epsilon)$ contains a $k$-dimensional linear subspace modulo $\ker L_{j}$.
-An analogous statement holds for $L_{j}^{\mathrm{cone}}$.
-Proof.
-Direct application of Min-Max theorem [1]. ∎
-Interpretation.
-Eigenvalues represent the critical slack levels at which new feasible directions appear.
-Small eigenvalues correspond to weakly violated constraints, while large eigenvalues correspond to strongly violated constraints.
+The parameter $\delta$ represents a budget of inconsistency.
+Increasing $\delta$ allows progressively larger violations of consistency or grounding constraints, yielding a persistence-style notion of feasibility.
 5.3 Integrated and Local Spectral Witnesses
 5.3.1 Global Spectral Witness
-Rather than relying on a single eigenvalue, we aggregate information across the feasibility filtration.
+To summarize the spectral filtration by scalar diagnostics, we introduce global spectral witnesses that aggregate low-energy spectral contributions across all feasible modes, quantifying the magnitude and persistence of inconsistency at a given cochain degree.
+Let $\{\lambda_{\ell}^{(j)}\}_{\ell\geq 1}$ denote the nonzero eigenvalues of the $j$-th sheaf (or cone) Laplacian $L_{j}$, counted with multiplicity.
+Let $w:\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$ be a fixed weighting function.
+Typical choices include
+$w_{\mathrm{unif}}(\lambda)=1,\qquad w_{\mathrm{inv}}(\lambda)=\lambda^{-1},\qquad w_{\mathrm{heat}}(\lambda)=e^{-t\lambda}.$
 Definition 11 (global spectral witness).
-Fix $0\leq\epsilon_{0}<\epsilon_{1}$. For $j\in\{0,1\}$, define (global) spectral witness at degree $j$ by
-$\mathcal{I}_{j}\;:=\;\int_{\epsilon_{0}}^{\epsilon_{1}}\Big(\min_{\begin{subarray}{c}x\neq 0\\ x\perp\ker L_{j}\end{subarray}}\frac{E_{j}(x)}{\|x\|^{2}}\;\wedge\;\epsilon\Big)\,d\epsilon,$
-and analogously for the cone Laplacian, (global) coned spectral witness at degree $j$ is
-$\mathcal{I}_{j}^{\mathrm{cone}}\;:=\;\int_{\epsilon_{0}}^{\epsilon_{1}}\Big(\min_{\begin{subarray}{c}x\neq 0\\ x\perp\ker L_{j}^{\mathrm{cone}}\end{subarray}}\frac{E_{j}^{\mathrm{cone}}(x)}{\|x\|^{2}}\;\wedge\;\epsilon\Big)\,d\epsilon,$
-where $a\wedge b:=\min\{a,b\}$.
+Fix $0\leq\delta_{0}<\delta_{1}$ and a nonnegative weighting function $w:\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$.
+Let $\{\lambda^{(j)}_{\ell}\}_{\ell\geq 1}$ denote the nonzero eigenvalues of the $j$-th (regular or cone) Laplacian, counted with multiplicity.
+We define the multi-mode global spectral witness by
+$\mathcal{I}_{j}(\delta_{0},\delta_{1})\;:=\;\sum_{\lambda^{(j)}_{\ell}\in(0,\delta_{1}]}(\delta_{1}-\max\{\delta_{0},\lambda^{(j)}_{\ell}\})\,w(\lambda^{(j)}_{\ell}).$
+Special case: gap-based witness.
+If the weighting function is chosen as $w(\lambda)=\mathbf{1}_{\{\lambda=\lambda_{\min}\}}$, then the global spectral witness reduces to a function of the smallest nonzero eigenvalue alone.
+This recovers the gap-based global witness used in earlier versions of this work, now interpreted as a single-mode special case of the general formulation.
+In the diagnostic experiments reported in Section 7, this gap-based global spectral witness is used for simplicity and numerical stability; the general multi-mode definition is included to emphasize that the framework itself does not depend on privileging a single spectral mode.
 Interpretation.
-The quantity $\mathcal{I}_{j}$ measures how long infeasibility persists under increasing slack at degree $j$.
-The cone version $\mathcal{I}_{j}^{\mathrm{cone}}$ measures persistence of infeasibility once ambient grounding is enforced.
-Integrating over $\epsilon$ yields a persistence-style invariant without altering the base topology.
-Remark 9 (Role of the kernel and spectral gaps).
-The dimension of $\ker L_{j}$ encodes the exact cohomology group $H^{j}(K;\mathcal{F})$.
-Accordingly, all spectral witnesses are defined on the orthogonal complement of the kernel in order to factor out gauge symmetries associated with exact consistency.
-As a consequence, spectral witnesses measure deviation from exact consistency rather than consistency itself.
-In particular, a small spectral witness should not be interpreted as evidence of consistency, but rather as indicating that any inconsistency present is weak or poorly resolved at the given spectral scale.
-Summary
-The base Laplacians $(L_{0},L_{1})$ quantify ungrounded consistency and obstruction, while the cone Laplacians $(L_{0}^{\mathrm{cone}},L_{1}^{\mathrm{cone}})$ quantify grounded feasibility.
-The $\epsilon$-feasible sets induce a spectral filtration whose critical values are eigenvalues, and the resulting spectral witnesses provide a quantitative measure of counterfactual inconsistency.
+Each eigenmode contributes to the global spectral witness in proportion to the range of spectral slack over which it remains feasible.
+Modes with smaller eigenvalues persist longer as $\delta$ increases and therefore contribute more heavily to the witness.
+Under this formulation, the global spectral witness measures the total low-energy spectral mass associated with inconsistency, without privileging any single extremal eigenvalue.
 5.3.2 Local spectral witnesses
-Global spectral witnesses certify the existence and persistence of inconsistency, but they do not identify its spatial origin.
-To localize inconsistency to specific cells, we refine the $\epsilon$-feasible spectral filtration by decomposing consistency energy into cellwise contributions associated with low-energy spectral modes.
+Global spectral witnesses summarize how inconsistency accumulates across the spectral filtration, but they do not resolve where such inconsistency is supported spatially on the cell complex.
+To obtain cell-level diagnostics, we refine the same low-energy spectral structure by decomposing consistency energy across individual cells.
+Local spectral witnesses assign to each cell a weighted contribution from low-energy spectral modes, yielding a spatial attribution of inconsistency that is consistent with the global spectral witness and evolves coherently with the spectral threshold.
 Definition 12.
@@
 $L_{j}v_{\ell}=\lambda_{\ell}v_{\ell},\qquad 0=\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\cdots.$
-Fix $\epsilon>0$ and define the spectral feasible subspace
-$\mathcal{V}_{j}(\epsilon)\;:=\;\mathrm{span}\{\,v_{\ell}\;:\;\lambda_{\ell}\leq\epsilon\,\}\;\subseteq\;C^{j}(K;\mathcal{F}),$
+Fix $\delta>0$ and define the spectral feasible subspace
+$\mathcal{V}_{j}(\delta)\;:=\;\mathrm{span}\{\,v_{\ell}\;:\;\lambda_{\ell}\leq\delta\,\}\;\subseteq\;C^{j}(K;\mathcal{F}),$
 with the understanding that the kernel $\ker L_{j}$ is excluded when $j>0$.
@@
 $L_{j}v_{\ell}=\lambda_{\ell}v_{\ell},\qquad\lambda_{\ell}\geq 0,$
-and let $w:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ be a nonnegative weighting function supported on a low-energy regime. For a spectral threshold $\epsilon>0$, we define the local spectral witness
-$W_{j,\epsilon}:K_{j}\longrightarrow\mathbb{R}_{\geq 0}$
+and let $w:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ be a nonnegative weighting function supported on a low-energy regime. For a spectral threshold $\delta>0$, we define the local spectral witness
+$W_{j,\delta}:K_{j}\longrightarrow\mathbb{R}_{\geq 0}$
 by
-$W_{j,\epsilon}(e):=\sum_{\lambda_{\ell}\leq\epsilon}w(\lambda_{\ell})\left(\sum_{\begin{subarray}{c}c_{j+1}\in K_{j+1}\\ c_{j+1}\geq e\end{subarray}}\bigl\|(d_{j}v_{\ell})[c_{j+1}]\bigr\|^{2}\;+\;\sum_{\begin{subarray}{c}c_{j-1}\in K_{j-1}\\ c_{j-1}\leq e\end{subarray}}\bigl\|(d_{j-1}^{T}v_{\ell})[c_{j-1}]\bigr\|^{2}\right).$
+$W_{j,\delta}(e):=\sum_{\begin{subarray}{c}\lambda_{\ell}\leq\delta\\ v_{\ell}\in\mathcal{V}_{j}(\delta)\end{subarray}}w(\lambda_{\ell})\left(\sum_{\begin{subarray}{c}e\prec c_{j+1}\in K_{j+1}\end{subarray}}\bigl\|(d_{j}v_{\ell})[c_{j+1}]\bigr\|^{2}\;+\;\sum_{\begin{subarray}{c}e\succ c_{j-1}\in K_{j-1}\end{subarray}}\bigl\|(d_{j-1}^{T}v_{\ell})[c_{j-1}]\bigr\|^{2}\right).$
 Here $(d_{j}v_{\ell})[c_{j+1}]\in\mathcal{F}(c_{j+1})$ and $(d_{j-1}^{T}v_{\ell})[c_{j-1}]\in\mathcal{F}(c_{j-1})$ denote the components of $d_{j}v_{\ell}\in C^{j+1}(K;\mathcal{F})$ and $d_{j-1}^{T}v_{\ell}\in C^{j-1}(K;\mathcal{F})$, respectively, under the canonical direct-sum decompositions of the cochain spaces.
-An analogous definition applies to the cone Laplacian $L_{j}^{\mathrm{cone}}$, yielding $\mathcal{W}_{j,\epsilon}^{\mathrm{cone}}(c)$.
+An analogous definition applies to the cone Laplacian $L_{j}^{\mathrm{cone}}$, yielding $\mathcal{W}_{j,\delta}^{\mathrm{cone}}(c)$.
 Remark 9.
@@
 Interpretation
-  • For $j=0$, $\mathcal{W}_{0,\epsilon}(c)$ localizes inconsistency to edges (and, by aggregation, to nodes), measuring which local constraints are violated by low-energy modes.
-  • For $j=1$, $\mathcal{W}_{1,\epsilon}(c)$ localizes obstruction both to (j+1)(j{+}1)-cells (triangle-level twists via $d_{1}$) and to edges (divergence or gauge components via $d_{0}^{\top}$).
-  • Cone versions $\mathcal{W}_{j,\epsilon}^{\mathrm{cone}}$ localize inconsistency that persists under ambient grounding, isolating defects that are invisible to ungrounded analysis.
-As $\epsilon$ increases, additional spectral modes contribute, yielding a persistence-style evolution of localized inconsistency.
-In this sense, $\mathcal{W}_{j,\epsilon}$ provides a cell-level refinement of the global spectral witness, identifying not only whether inconsistency exists, but where it is concentrated.
+  • For $j=0$, $\mathcal{W}_{0,\delta}(c)$ localizes inconsistency to edges (and, by aggregation, to nodes), measuring which local constraints are violated by low-energy modes.
+  • For $j=1$, $\mathcal{W}_{1,\delta}(c)$ localizes obstruction both to (j+1)(j{+}1)-cells (triangle-level twists via $d_{1}$) and to edges (divergence or gauge components via $d_{0}^{\top}$).
+  • Cone versions $\mathcal{W}_{j,\delta}^{\mathrm{cone}}$ localize inconsistency that persists under ambient grounding, isolating defects that are invisible to ungrounded analysis.
+As $\delta$ increases, additional spectral modes contribute, yielding a persistence-style evolution of localized inconsistency.
+In this sense, $\mathcal{W}_{j,\delta}$ provides a cell-level refinement of the global spectral witness, identifying not only whether inconsistency exists, but where it is concentrated.
 6 Taxonomy of Inconsistency Mechanisms
@@
 Table 3 compares a hidden twist and a noisy trivial bundle.
-Both have trivial kernel, but differ in their smallest nonzero eigenvalue and in the integrated spectral witness $\mathcal{I}_{0}$.
+Both have trivial kernel, but differ in their smallest nonzero eigenvalue and in the integrated (gap-based) spectral witness $\mathcal{I}_{0}$.
 These quantities measure the magnitude of inconsistency, i.e.
@@
 Figure 1: Local spectral witnesses for hidden twist (top row) and noisy trivial assignment (bottom row).
-Each row visualizes local spectral witnesses $\mathcal{W}_{j,\epsilon}$ on the same cycle complex, comparing intrinsic (base) Laplacian channels with grounded (mapping-cone) diagnostics.
+Each row visualizes local spectral witnesses $\mathcal{W}_{j,\delta}$ on the same cycle complex, comparing intrinsic (base) Laplacian channels with grounded (mapping-cone) diagnostics.
 Shown are heatmaps induced by low-energy eigenmodes of the intrinsic $j=0$ and $j=1$ Laplacians, together with the relative cone channel corresponding to the intrinsic $j=1$ structure.
@@
 This step extracts an intrinsic local representation while discarding redundant or degenerate directions.
+Remark 10.
+Although node-level features are used throughout the experiments in this paper, the framework does not assume that the hypothetical structure is vertex-generated.
 B.2 Clique Complex Construction
@@
 Rather, it serves as a discrete indicator of exact consistency, while the spectral gap on its orthogonal complement provides a robust quantitative witness of inconsistency.
+B.8 Hypothetical structures and stability across admissible sheaf realizations
+We formalize the sense in which the proposed spectral diagnostics are independent of specific sheaf-construction choices, while remaining sensitive to the underlying structural hypotheses.
+Hypothetical structure.
+A hypothetical structure $\mathcal{H}$ consists of partial cellular data specified on a cell complex $K$, including:
+  • a choice of cells $\Sigma\subseteq K$ on which stalk data are prescribed;
+  • for each $\sigma\in\Sigma$, a finite-dimensional inner product space $\mathcal{H}(\sigma)$ representing structural features at $\sigma$;
+  • optionally, partial compatibility data or constraints between prescribed stalks.
+No assumption is made that $\Sigma$ consists only of vertices, nor that restriction maps are fully specified. In particular, $\mathcal{H}$ need not define a sheaf.
+Family of constructions.
+Given a hypothetical structure $\mathcal{H}$, we consider a family of admissible construction procedures indexed by parameters $\theta\in\Theta$. Each construction specifies:
+  • how to extend stalk assignments from $\Sigma$ to additional cells when required;
+  • how to choose bases or coordinate representations for stalks;
+  • how to deduce or approximate restriction maps between cells;
+  • how to select auxiliary parameters such as tolerances or truncation thresholds.
+These choices are algorithmic and modeling-dependent, and are not regarded as intrinsic to $\mathcal{H}$.
+Realized sheaves and admissible objects.
+Each construction $\theta\in\Theta$ produces a cellular sheaf
+$\mathcal{F}_{\theta}\;=\;\mathcal{C}_{\theta}(\mathcal{H})$
+on $K$, equipped with inner product stalks and linear restriction maps.
+Different choices of $\theta$ may yield non-isomorphic sheaves, supported on different collections of cells or with different internal dimensions.
+We denote by $\mathsf{Shv}(\mathcal{H})$ the resulting class of admissible realizations of $\mathcal{H}$, i.e.
+the collection of cellular sheaves arising from constructions compatible with the same hypothetical structure.
+No canonical morphisms between elements of $\mathsf{Shv}(\mathcal{H})$ are assumed to exist.
+Spectral indicators.
+Let $\mathcal{D}$ denote any spectral diagnostic derived from the sheaf Laplacians (or cone Laplacians), such as spectral gaps, integrated spectral witnesses, or localized energy profiles.
+The framework does not assert that $\mathcal{D}(\mathcal{F}_{\theta})$ is invariant under changes of $\theta$.
+Rather, it asserts that for admissible constructions $\theta,\theta^{\prime}\in\Theta$, the quantities
+$\mathcal{D}(\mathcal{F}_{\theta})\quad\text{and}\quad\mathcal{D}(\mathcal{F}_{\theta^{\prime}})$
+are quantitatively comparable and stable up to explicitly controlled scale parameters.
+Interpretation.
+In this sense, the spectral diagnostics are understood as properties of the underlying hypothetical structure $\mathcal{H}$, as revealed through a class of admissible sheaf realizations, rather than invariants of any single constructed sheaf.
+We refer to this principle as stability across admissible sheaf realizations.